### v2

# via Tweedie Discrete Diffusion

Abstract
Vision–Language–Action (VLA) models offer a unified framework for robotic manipulation by integrating visual perception, language understanding, and control generation. However, existing VLA systems still struggle to generalize across diverse tasks, scenes, and camera viewpoints, and often produce coarse or unstable actions. We argue that these limitations are closely tied to the structural properties of actions in VLA settings, including the inherent multi-peaked nature of action distributions, the token-based symbolic reasoning of pretrained VLM/VLA backbones, and the effective finite resolution imposed by real-world robotic control. Motivated by these properties, we introduce $\mathcal{E}_{0}$, a tweedie discrete diffusion framework that formulates action generation as iterative denoising over quantized action tokens. By operating in a discrete action space with a principled diffusion process, $\mathcal{E}_{0}$ naturally aligns with token-based reasoning, supports fine-grained yet executable action control, and avoids the distributional mismatch of masking-based discrete diffusion. We further introduce a spherical viewpoint perturbation augmentation to enhance robustness to camera shifts without additional data. Experiments on LIBERO, VLABench, ManiSkill, and a real-world Franka arm demonstrate that $\mathcal{E}_{0}$ achieves state-of-the-art performance across 14 diverse environments, outperforming strong baselines by 10.7% on average.
Vision-Language-Action Models, Embodied AI
1 Introduction
Figure 1: Overview of action modeling paradigms. (a) Discrete modeling: Traditional autoregressive (AR) approaches (Brohan et al., 2022; Zitkovich et al., 2023; Kim et al., 2024) and recent mask-based discrete diffusion methods (Liang et al., 2025), which operate over a small discrete action vocabulary. (b) Continuous modeling: Continuous diffusion–based policies (Liu et al., 2024; Xu et al., 2025) and AR–diffusion hybrids (Black et al., 2024; Intelligence et al., 2025; Lin et al., 2025) that regress continuous actions. (c) Our approach: $\mathcal{E}_{0}$ integrates AR-style conditioning with a tweedie discrete diffusion framework, enabling efficient action generation while preserving compatibility with pretrained vision–language backbones and supporting fine-grained action control.
Robotic manipulation in open-ended environments requires models that can perceive complex visual scenes, follow natural language instructions, and generate precise, reliable actions. Vision–Language–Action (VLA) models have recently emerged as a unified paradigm for this goal by integrating visual perception, language understanding, and control generation within a single framework. Leveraging large-scale multimodal pretraining, these models aim to generalize across tasks, scenes, and object categories. However, despite their promise, existing VLA systems still struggle to cope with diverse task instructions, environment configurations, and camera viewpoints, and often produce coarse or unstable actions that fail in fine-grained manipulation.
Recent VLA systems have explored a diverse range of formulations for action generation. Representative approaches include autoregressive token prediction over discretized action spaces (Brohan et al., 2022; Zitkovich et al., 2023; Kim et al., 2024; Pertsch et al., 2025), masked diffusion over discrete action tokens (Nie et al., 2025; Wu et al., 2025; Ye et al., ), as well as continuous diffusion models that iteratively denoise real-valued action trajectories (Black et al., 2024; Intelligence et al., 2025; Lin et al., 2025). While differing in architectural design and training objectives, these methods largely adopt action representations inherited from existing sequence modeling or diffusion paradigms, rather than being derived from a principled analysis of action features specific to VLA settings.
To better understand the challenges of action modeling in VLA systems, it is useful to examine the structural properties of actions in this setting. First, robotic actions are inherently multi-peaked (Florence and Lynch, 2021; Shafiullah et al., 2022; Florence et al., 2022): under the same task specification and visual context, multiple distinct action sequences may successfully accomplish the task, making action prediction fundamentally a distribution modeling problem rather than single-valued regression (see Appendix K.1 for additional analysis). Second, action generation in VLA systems is tightly coupled with pretrained vision–language backbones, where high-level reasoning, instruction following, and semantic conditioning are mediated through discrete symbolic tokens (Liang et al., 2025; Li et al., 2025; Luo et al., 2026). These backbones are optimized with token-level classification objectives, under which semantic information is primarily encoded through relative feature directions, shaping the representational and optimization characteristics of VLM/VLA models (see Appendix K.2 for additional analysis). Third, although physical execution evolves in continuous time, practical robotic systems operate under finite control rates, encoder resolution, actuation latency, and environmental noise. These factors collectively impose an effective finite resolution on action execution at the policy level, which may differ substantially from nominal hardware specifications due to real-world physical constraints (see Appendix K.4 for additional analysis). Taken together, these properties highlight fundamental constraints on action modeling in VLA systems, motivating formulations that can capture multi-peaked behaviors, align with token-based reasoning, and operate under finite-resolution execution limits.
For multi-peaked action distributions, discrete representations naturally preserve multiple valid modes without collapsing them into a single averaged prediction. For action generation mediated by discrete symbolic tokens in pretrained vision–language backbones, discrete action spaces enable representational and optimization-level alignment with token-based reasoning. Finally, the effective finite resolution imposed by real-world robotic control makes discrete abstractions well matched to the granularity at which actions can be reliably generated and executed. Collectively, these observations suggest that discreteness in action modeling follows naturally from the multi-peaked structure of actions, the token-based reasoning of VLA models, and the finite resolution of real-world control.
To this end, we propose $\mathcal{E}_{0}$, a tweedie discrete diffusion framework that formulates action generation as iterative denoising over Gaussian-noised one-hot action vectors (see Fig. 1(c)). By applying Gaussian noise directly to discrete action spaces, $\mathcal{E}_{0}$ follows Tweedie’s formulation and maintains forward–reverse consistency, avoiding the distributional mismatch inherent in mask-based diffusion (see Appendix L for additional analysis). Unlike AR-based methods, our approach supports arbitrarily fine discretization bins, enabling high-resolution action modeling beyond the constraints of language tokenizers. At the same time, by operating in a discrete space, $\mathcal{E}_{0}$ remains aligned with the symbolic representations used in pretrained VLM/VLA backbones, strengthening semantic conditioning. Furthermore, because discrete diffusion models the true quantized nature of robot control, its Bayes-optimal denoiser captures the correct discrete action distribution and yields stronger generalization than continuous models.
To further enhance robustness, we introduce a spherical viewpoint perturbation augmentation that perturbs input observations based on the relative camera–scene geometry. This improves cross-view consistency and reduces sensitivity to camera shifts without requiring additional data collection, addressing a long-standing practical challenge.
In summary, our contributions are threefold:
  1. 1.
    We propose $\mathcal{E}_{0}$, a tweedie discrete diffusion framework for action modeling that supports arbitrarily fine discretization. This formulation enables high-precision action representation while maintaining efficient inference, and preserves compatibility with pretrained vision–language models for accurate perception, grounding, and manipulation.
  2. 2.
    We introduce a spherical viewpoint perturbation augmentation together with a relative spherical embedding mechanism. This explicitly models dynamic camera perturbations and significantly improves cross-view consistency and robustness in action generation.
  3. 3.
    We demonstrate the effectiveness of $\mathcal{E}_{0}$ through extensive experiments across multiple simulation benchmarks and real-world Franka manipulation tasks, covering a wide range of scenes, object types, and task complexities.
2 Related Work
Figure 2: Overview and detailed illustration of $\mathcal{E}_{0}$. (a) Overall architecture of the proposed model. (b) Training and inference pipeline, showing how inputs are encoded, diffused, and decoded into executable action sequences.
Autoregressive VLA Models. Early efforts such as RT-1 (Brohan et al., 2022) and RT-2 (Zitkovich et al., 2023) pioneered VLA models, demonstrating the potential of scaling language-conditioned policies. Building on this foundation, OpenVLA (Kim et al., 2024) released the first open-source autoregressive VLA by integrating Llama-2 (Touvron et al., 2023) with strong vision encoders (DINOv2 (Oquab et al., 2023), SigLIP (Zhai et al., 2023)), while SpatialVLA (Qu et al., 2025) introduced Ego3D position encoding to inject explicit 3D spatial cues. To enhance tokenized action generation, OpenVLA-OFT (Kim et al., 2025) proposed an optimized fine-tuning recipe with parallel decoding and continuous action representations. $\pi_{0}$ FAST (Pertsch et al., 2025) designed a frequency-aware tokenization scheme for efficient autoregressive training. In parallel, reasoning-augmented approaches such as CoT-VLA (Zhao et al., 2025) and generative rollouts (GR-1 (Wu et al., 2023), GR-2 (Cheang et al., 2024)) extend VLAs with textual or visual chain-of-thought to guide planning. Our work builds on these advances by preserving the strong contextual alignment afforded by autoregressive decoding, while addressing its inherent limitations in producing fine-grained continuous actions.
Diffusion-based VLA Models. The application of diffusion to robotic imitation learning was first introduced by Diffusion Policy (Chi et al., 2023), demonstrating its promise for continuous action generation. Subsequent works have advanced this paradigm through increasingly sophisticated designs: CogACT (Li et al., 2024) developed diffusion action transformers with favorable scaling properties, while RDT (Liu et al., 2024) exploited DiT’s cross-modal fusion to generate complete action sequences directly from noisy action chunks, later extended by A0 (Xu et al., 2025) with affordance-based reasoning. Generalist frameworks such as $\pi_{0}$ (Black et al., 2024) and $\pi_{0.5}$ (Intelligence et al., 2025) proposed powerful flow models for broad robotic control, achieving strong performance across diverse real-world tasks. Other recent directions include OneTwoVLA (Lin et al., 2025), which enables dynamic switching between reasoning and execution, GR00t N1 (Bjorck et al., 2025) with a dual-system design, and GO-1 (Bu et al., 2025), a scalable generalist policy using latent actions. Finally, Hybrid VLA (Liu et al., 2025) combines autoregressive and diffusion decoding to leverage their complementary strengths. In contrast, our work preserves the expressive continuous modeling ability of diffusion while maintaining the emergent language grounding and multimodal reasoning capacities of AR VLAs.
Discrete Diffusion. Recent advances have sparked growing interest in discrete diffusion models. A common formulation treats discrete diffusion as a masked modeling problem, in which noise is simulated through random token masking and the model is trained to reconstruct the original input (Nie et al., 2025; Wu et al., 2025; Ye et al., ). Recently, an attempt has been made to extend discrete diffusion to VLA models (Liang et al., 2025); however, their design follows a BERT-like (Devlin et al., 2019) masking mechanism and requires additional architectural complexity to achieve competitive performance. In contrast, our discrete diffusion provides a reformulation that operates directly on float-encoded one-hot class representations, avoiding both continuous latent diffusion and heuristic masking strategies. During inference, the model predicts a clean categorical state at each step by applying an $\arg\max$ operation over output probabilities and converting the result back into a one-hot vector. Such an autoregressive-style feedback mechanism enables gradual refinement of predictions. Building upon this foundation, our work makes more stable action generation and achieves competitive performance in complex control scenarios.
3 Method
3.1 Preliminaries
Continuous-Time Flow Matching. Flow matching (Lipman et al., 2022; Tong et al., 2023) formulates generative modeling as learning a continuous-time velocity field that transports samples from a simple noise distribution to a target data distribution. Given a clean sample $x_{0}$ and noise $\epsilon\sim\mathcal{N}(0,I)$, an intermediate point is constructed as
$x_{t}=t\cdot\epsilon+(1-t)\cdot x_{0},\quad t\in[0,1].$ (1)
The model is trained to predict the velocity
$u_{t}=\epsilon-x_{0},$ (2)
using a regression objective.
Inference via ODE Integration. At inference time, samples are generated by solving the ordinary differential equation
$\frac{dx_{t}}{dt}=v_{\theta}(x_{t},t),$ (3)
from $t=1$ to $t=0$ using a numerical solver. This process is deterministic and does not require an explicit noise variance schedule.
3.2 Architecture and Pipeline
VLM Backbone and Action Expert. As shown in Fig. 2(a), we build on PaliGemma (Beyer et al., 2024), an open-source VLM, and additionally incorporate a 300M-parameter action expert as the representation backbone. Unlike traditional diffusion-based approaches (Black et al., 2024; Intelligence et al., 2025) that map the outputs of a VLM into continuous action representations via an MLP, our method discretizes actions into bins with arbitrary granularity, while recording the corresponding statistics to enable accurate recovery of the original continuous actions. Notably, autoregressive VLAs (Brohan et al., 2022; Zitkovich et al., 2023; Kim et al., 2024) typically encode discrete actions by occupying tokens in the original vocabulary and fix the size of 256. Therefore, increasing the action resolution requires expanding the vocabulary, which incurs additional inference overhead and often necessitates retraining the model. In contrast, our method does not rely on the underlying token vocabulary: action discretization is implemented as a numerical mapping, enabling arbitrarily fine action granularity without increasing inference cost or modifying the model architecture.
Consistent with prior work (Brohan et al., 2022; Zitkovich et al., 2023; Kim et al., 2024; Liang et al., 2025), we adopt a quantile-based discretization scheme (using the 1st–$M$th percentiles for each dimension to mitigate the influence of outliers). This effectively filters out extreme values and ensures smooth and stable performance during robotic inference. A single-timestep action can be represented as $D_{a}=7$ (3 for translation, 3 for rotation, and 1 for the gripper). Alternatively, it can also be defined as $D_{a}=8$ ( 7 for joint angles and 1 for the gripper, e.g., Franka Research 3). In our experiments, we set the number of bins to $2048$. For action chunking, tokens from $H$ (with $H=50$ in our experiments) future timesteps are arranged into a fixed layout, yielding a total of $L_{a}=H\times D_{a}$ action positions.
Figure 3: Benchmarks for evaluation. (a) LIBERO (Liu et al., 2023): tasks with varying objects, layouts, and goals, including long-horizon settings. (b) ManiSkill (Tao et al., 2024): diverse fine-grained manipulation skills (push, pick, stack, insert, plug). (c) VLABench (Zhang et al., 2024): open-ended tasks requiring language grounding and commonsense reasoning (select toy/fruit/painting/poker/mahjong).
Training Pipeline. As shown in Fig. 2(b), our objective is to model the conditional distribution $p(A_{t}\mid o_{t})$, where $A_{t}=[a_{t},a_{t+1},\ldots,a_{t+H-1}]$ denotes an action chunk of horizon $H$ capturing temporal dependencies, and $o_{t}=[I_{t}^{1},\ldots,I_{t}^{n},\,l_{t},\,q_{t}]$ represents the multimodal observation at time $t$. Here, $I_{t}^{i}$ is the $i$-th RGB image among $n$ camera views, $l_{t}$ encodes the language instruction, and $q_{t}$ denotes the robot’s proprioceptive state (e.g., joint angles). This formulation enables the model to jointly reason over heterogeneous modalities and predict coherent future actions conditioned on rich contextual cues.
Visual features $I_{t}^{i}$ and proprioceptive states $q_{t}$ are processed by their respective encoders and linearly projected into the same embedding space as the language tokens, yielding a unified cross-modal representation. Each action $a_{t}$ is discretized using the 1st–$M$th percentiles, producing a sequence of discrete tokens $\tilde{A}_{t}$ of length $L$, represented as one-hot vectors.
During training, a timestep $\tau\in[0,1]$ is sampled, and Gaussian noise $\varepsilon\!\sim\!\mathcal{N}(0,I)$ is added to the discretized actions:
$\tilde{A}_{t}^{\tau}=\tau\varepsilon+(1-\tau)\tilde{A}_{t}.$ (4)
Since the discretized one-hot actions are inherently sharp, we apply an additional smoothing factor $\alpha=0.1$ on $\tilde{A}_{t}$ before adding noise, which improves training stability. Following (Black et al., 2024), we draw $\tau$ from a beta distribution instead of the uniform schedule (Lipman et al., 2022; Liu, 2022). This bias toward smaller $\tau$ (i.e., higher-noise regions) reflects the stronger multimodal constraints in embodied action prediction compared to image generation, encouraging the model to learn robust denoising behavior under high uncertainty.
Given the noisy action representation $\tilde{A}_{t}^{\tau}$ and observation $o_{t}$, the network outputs logits $v_{\theta}(\tilde{A}_{t}^{\tau},o_{t})$ defining a categorical distribution:
$p_{\theta}(A_{t}\mid\tilde{A}_{t}^{\tau},o_{t})=\mathrm{Softmax}\!\left(v_{\theta}(\tilde{A}_{t}^{\tau},o_{t})\right).$ (5)
Training minimizes the cross-entropy loss between predicted and ground-truth tokens:
$\mathcal{L}_{\mathrm{CE}}(\theta)=-\,\mathbb{E}_{t}\!\left[\sum_{h=1}^{H}\log p_{\theta}\!\left(A_{t,h}=\tilde{A}_{t,h}\mid\tilde{A}_{t}^{\tau},o_{t}\right)\right],$ (6)
which encourages accurate and temporally consistent sequence generation across the action horizon.
Inference Pipeline. During inference, we initialize the action sequence from Gaussian noise and perform a multi-step iterative refinement process. At the first iteration, the observation $o_{t}$ is encoded once by the model to construct a key–value cache $\mathit{KV}(o_{t})$ for cross-attention. This cache is reused across all subsequent refinement steps, such that only the action tokens are recomputed.
At each iteration $i\in\{1,\ldots,N\}$ (with $N=10$ in our experiments), the model takes as input the current noisy action representation $\tilde{A}_{t}^{\tau_{i}}$ together with the fixed observation cache $\mathit{KV}(o_{t})$, and predicts a categorical distribution over discrete action tokens. The predicted tokens are converted into one-hot representations, which are used to construct an updated action vector for the next iteration. Noise annealing is implicitly controlled by the scalar timestep $\tau_{i}$, which monotonically decreases from $1$ to 0 over refinement steps. Specifically, the next noisy action representation is obtained as:
$\tilde{A}_{t}^{\tau_{i+1}}=\tau_{i+1}\,\varepsilon+(1-\tau_{i+1})\,\hat{A}_{t}^{(i)},\quad\varepsilon\sim\mathcal{N}(0,I),$ (7)
where $\varepsilon$ is fixed-variance isotropic Gaussian noise and $\hat{A}_{t}^{(i)}$ denotes the discrete action estimate at iteration $i$.
After $N$ refinement steps, the final discrete action tokens are deterministically detokenized into continuous actions, yielding the reconstructed action chunk $A_{t}=[a_{t,1},\ldots,a_{t,H}]$. Further details are provided in Appendix A.1.
3.3 Spherical Perspective Generalization
To improve robustness against dynamic camera perturbations, we introduce a spherical warping augmentation paired with a relative spherical embedding. Given an RGB image and camera intrinsics, we unproject each pixel $(u,v)$ to a 3D point at a fixed depth $d_{0}$, apply a yaw–pitch rotation $R(\Delta\phi,\Delta\theta)$, and reproject to obtain a warped image:
$u^{\prime},v^{\prime}]=\Pi\!\big(R(\Delta\phi,\Delta\theta)\,\Pi^{-1}(u,v,d_{0})\big),$ (8)
which simulates camera motion on a viewing sphere. In parallel, each view is associated with a 3D offset $\delta=(d,\theta,\phi)$ capturing radial, horizontal, and vertical displacements. A learnable projection $f_{\mathrm{proj}}$ maps this offset into the token space and is added to image tokens:
$z^{\prime}_{\mathrm{img}}=z_{\mathrm{img}}+f_{\mathrm{proj}}(\delta).$ (9)
Joint training with warped images and relative embeddings reduces reliance on fixed viewpoints and improves cross-view consistency in action generation. Further details are provided in the section A.2 of the appendix.
4 Experiment
Table 1: Performance comparison across multiple benchmarks (LIBERO, VLABench, ManiSkill). Each benchmark contains several task subsets, and the final column reports the average success rate (%). A value of 0 indicates the model completely failed on this task, while “–” denotes cases where evaluation was not conducted due to model or dataset constraints. A “–” in the final average column means the model was not tested on all benchmarks (direct averaging would be misleading). Bold numbers mark the best performance per column.
Model LIBERO (Liu et al., 2023) VLABench (Zhang et al., 2024) ManiSkill (Tao et al., 2024) Avg.
Spatial Object Goal Long Toy Fruit Painting Poker Mahjong Insert Pick Stack Plug Push
Diffusion Policy (Chi et al., 2023) 78.3 92.5 68.3 50.5 - - - - - 0.0 40.0 80.0 0.0 88.0 -
MDT (Reuss et al., 2024) 78.5 87.5 73.5 64.8 - - - - - - - - - - -
RDT (Liu et al., 2024) - - - - - - - - - 13.2 77.2 74.0 1.2 100.0 -
Dita (Hou et al., 2025) 84.2 96.3 85.4 63.8 - - - - - - - - - - -
TraceVLA (Zheng et al., 2024) 84.6 85.2 75.1 54.1 - - - - - - - - - - -
SpatialVLA (Qu et al., 2025) 88.2 89.9 78.6 55.5 - - - - - - - - - - -
OpenVLA (Kim et al., 2024) 84.7 88.4 79.2 53.7 - - - - - 0.0 8.0 80.0 0.0 88.0 -
Octo (Mees et al., 2024) 78.9 85.7 84.6 51.1 - - - - - 0.0 0.0 0.0 0.0 0.0 -
$\pi_{0}$ FAST (Pertsch et al., 2025) 96.4 96.8 88.6 60.2 46.0 42.0 26.0 30.0 20.8 0.0 80.0 52.0 0.0 92.0 52.2
$\pi_{0}$ (Black et al., 2024) 96.8 98.8 95.8 85.2 54.0 48.0 16.0 6.0 7.0 4.0 60.0 48.0 0.0 100.0 51.4
$\pi_{0.5}$ (Intelligence et al., 2025) 95.4 98.4 97.2 89.6 24.0 18.0 36.0 20.0 6.5 8.0 56.0 56.0 4.0 92.0 50.1
$\mathcal{E}_{0}$ (ours) 97.2 99.4 95.0 92.2 54.0 34.0 12.0 72.0 18.8 24.0 76.0 72.0 4.0 100.0 60.8
4.1 Training Setting
The pretrained VLM is based on Gemma (Team et al., 2024), which is configured as $\emph{width}=2048$, $\emph{mlp\_dim}=16384$, $\emph{depth}=18$, $\emph{num\_heads}=8$, and $\emph{head\_dim}=256$. The action expert adopts the same architecture but with a smaller configuration: $\emph{width}=1024$, $\emph{mlp\_dim}=4096$, $\emph{depth}=18$, $\emph{num\_heads}=8$, and $\emph{head\_dim}=256$. All models were fine-tuned on their corresponding datasets for 24 hours using a single NVIDIA RTX RPO6000 GPU. Each model was trained for a total of 30,00030{,}000 steps with a batch size of $32$. The learning rate followed a cosine decay schedule, implemented as CosineDecaySchedule, with a warm-up phase of 1,0001{,}000 steps, a peak learning rate of $5\times 10^{-5}$ and a final learning rate of $5\times 10^{-6}$. The optimizer is AdamW (Adam et al., 2019), configured with a gradient clipping norm of $1.0$. An exponential moving average with a decay rate of $0.999$ was applied to stabilize training. During inference, a single NVIDIA RTX 3090 GPU was used for model evaluation and deployment on the server side.
While our model adopts the same backbone network as $\pi_{0}$ FAST (Pertsch et al., 2025), $\pi_{0}$ (Black et al., 2024), and $\pi_{0.5}$ (Intelligence et al., 2025), the action modeling in these methods relies on autoregressive or diffusion-based action heads. Consequently, we consider them as baselines and retrain them using the same experimental setup to ensure a fair comparison and to control for variations in training settings and environments.
4.2 Simulation Experiment
As shown in Fig. 3, we evaluate our proposed $\mathcal{E}_{0}$ on three representative simulation benchmarks. LIBERO (Liu et al., 2023) provides a suite of manipulation tasks designed for benchmarking robot learning. It consists of four categories—Libero-Spatial, Libero-Object, Libero-Goal, and Libero-10. Following RDT (Liu et al., 2024), we adopt a representative subset of five manipulation tasks with varying levels of difficulty—PegInsertionSide, PickCube, StackCube, PlugCharger, and PushCube—for comparative evaluation in ManiSkill (Tao et al., 2024). VLABench (Zhang et al., 2024) is specifically designed to evaluate general-purpose robotic systems built upon large multimodal models. It emphasizes language understanding, commonsense reasoning, and multi-step task planning, covering five dimensions: grid texture perception, spatial reasoning, commonsense and world knowledge, semantic understanding, and physical laws.
Analysis on LIBERO. This benchmark includes a wide range of models, many of which employ sophisticated training strategies and architectural enhancements to maximize performance. Despite the highly competitive nature of the benchmark, our model achieved the highest average performance (96%), demonstrating the effectiveness of the proposed approach even when trained with mixed data. To ensure fairness, we trained $\pi_{0}$, $\pi_{0}$ FAST, and $\pi_{0.5}$ using the same configuration and validated them under identical settings. Experimental results show that our model exhibits more accurate grasping and is better at executing task instructions in some of the more complex scenarios. Further details are provided in the section D of the appendix.
[…]
Analysis on VLABench. We observe that continuous action models exhibit noticeable performance degradation on tasks that demand explicit reasoning or fine-grained visual discrimination. This behavior is anticipated, as $\pi_{0}$ relies solely on action supervision during training, without incorporating explicit constraints for contextual reasoning. Consequently, its autoregressive backbone often struggles to maintain semantic coherence during inference. In contrast, $\pi_{0}$ FAST shows marked improvements through its autoregressive generation framework, which enhances task-level reasoning. However, it still suffers from relatively imprecise action generation. Our proposed method, $\mathcal{E}_{0}$, achieves a balanced trade-off between these paradigms. By integrating the contextual grounding capabilities of vision–language reasoning with the fine-grained control offered by discrete diffusion, it enables robust multimodal understanding and generates precise, semantically consistent action trajectories. As shown in Fig. 4, $\pi_{0}$ and $\pi_{0}$ FAST misinterpret the textual instruction and select incorrect suits, while $\pi_{0.5}$ identifies the correct card but fails to grasp it precisely. In contrast, $\mathcal{E}_{0}$ both recognizes the correct suit and executes an accurate grasp, demonstrating strong semantic understanding and precise control.
Figure 4: Comparison on the VLABench benchmark. In the task “pick up the spade 3”, our $\mathcal{E}_{0}$ correctly identifies and precisely grasps the target card, showing superior multimodal reasoning and control ability.
4.3 Real World Experiment
For real-world validation, we employ a Franka Research 3 robotic arm and evaluate our method across eight categories of manipulation tasks (see Fig. 5). We utilize Gello (Wu et al., 2024) to control the Franka arm and collect demonstration data. For the short-horizon tasks, we collect 50 trajectories per task, while for the long-horizon tasks, we collect 80 trajectories each. We train our model on a mixture of these trajectories and evaluate its performance by executing 20 trials per task. As shown in the Tab. 2, our model achieves the highest average success rate across all tasks. These results demonstrate $\mathcal{E}_{0}$’s capability to maintain semantic coherence and precise control across extended action sequences. Moreover, we observe impressive generalization to previously unseen configurations. For instance, in the stack block task, even when the positions of the red and green blocks are swapped—differing from the training data—the model accurately identifies their colors and completes the stacking correctly (see Fig. 5). Similarly, in the pick block twice task, while the green plate and the red and orange blocks were fixed during data collection, the model successfully handles varying placements during validation, demonstrating robust generalization and adaptability. Further details are provided in the section H of the appendix.
(a) Short-horizon tasks.
(b) Long-horizon tasks.
Figure 5: Performance on real-world robotic experiments. (a) Short-horizon tasks (press button, close door, pull drawer, pick block, stack block, stack block unseen). (b) Long-horizon tasks (pick block twice, pull drawer and put in block, and put in plate and close door).
Table 2: Task success rates in real-world robot experiments. We evaluate policy performance across both short-horizon and long-horizon manipulation tasks. Bold numbers indicate the best performance in each column.
Model Short-horizon Tasks Long-horizon Tasks Average
Pick Press Stack Pull Close Pick Twice Open & Put Put & Close
SR (%) $\uparrow$ SR (%) $\uparrow$ SR (%) $\uparrow$ SR (%) $\uparrow$ SR (%) $\uparrow$ SR (%) $\uparrow$ SR (%) $\uparrow$ SR (%) $\uparrow$ SR (%) $\uparrow$
$\pi_{0}$ (Black et al., 2024) 60.0 60.0 40.0 30.0 45.0 45.0 55.0 10.0 43.1
$\pi_{0}$ FAST (Pertsch et al., 2025) 20.0 20.0 5.0 10.0 15.0 10.0 0.0 0.0 10.0
$\mathcal{E}_{0}$ (ours) 75.0 70.0 50.0 35.0 35.0 60.0 40.0 20.0 45.6
Figure 6: Comprehensive ablation analysis of key hyperparameters in the LIBERO environments. We investigate four crucial design factors influencing our model: (a) Discretization bins—increasing bin granularity enhances precision up to 2048 bins, beyond which gains saturate; (b) Action horizon—a moderate prediction length balances reactivity and temporal consistency; (c) Action dimensions—embedding sizes slightly above the dataset’s action space yield the best expressiveness–robustness trade-off; and (d) One-hot smooth factor—moderate decay values smooth discrete logits, stabilizing diffusion and improving overall success rate.
Figure 7: Comparison on the VLABench Select Painting task. $\mathcal{E}_{0}$ shows more consistent visual–language grounding and smoother control than $\pi_{0.5}$, which often succeeds only by chance through default center-button presses.
4.4 Ablation Study
Robustness to camera viewpoint changes. To enhance policy robustness under dynamic camera perturbations, we introduce a spherical warping augmentation coupled with a relative spherical embedding. This design mitigates the common issue of viewpoint overfitting in visuomotor policies trained under fixed camera setups. By encoding 3D camera displacements into the latent representation, the model achieves consistent feature alignment across varying viewpoints.
As shown in Tab. 3, both $\pi_{0}$ and our $\mathcal{E}_{0}$ exhibit substantial performance degradation without view augmentation (average SR of 19.7% and 66.5%, respectively), indicating strong sensitivity to camera changes. Incorporating spherical view augmentation (+view) markedly improves robustness, boosting average SR to 50.8% and 83.9%. These results demonstrate that our view-aware design significantly enhances generalization to unseen or dynamic camera configurations, a key requirement for reliable real-world deployment.
Table 3: Evaluation under camera viewpoint perturbations in the LIBERO environments. We evaluate the robustness of different models when camera positions and orientations are dynamically perturbed.
Model Libero-Spatial Libero-Object Libero-Goal Libero-Long Average
SR (%) $\uparrow$ SR (%) $\uparrow$ SR (%) $\uparrow$ SR (%) $\uparrow$ SR (%) $\uparrow$
$\pi_{0}$ (Black et al., 2024) 28.0 16.6 31.6 2.6 19.7
$\pi_{0}$ (Black et al., 2024) + view 53.8 75.2 57.8 16.2 50.8 (+31.1)
$\mathcal{E}_{0}$ (Ours) 74.4 85.2 77.4 28.8 66.5
$\mathcal{E}_{0}$ (Ours) + view 92.0 96.8 83.0 63.8 83.9 (+22.6)
Critical hyperparameters in our model. We perform an extensive ablation study to examine the impact of key hyperparameters in the LIBERO environments. As shown in Fig. 6, finer action discretization improves control precision up to a moderate resolution, beyond which the enlarged vocabulary introduces optimization noise with limited gain. Matching the action dimension to the dataset’s intrinsic control size ensures representational completeness and stable diffusion, while excessive dimensionality harms smoothness. Moderate one-hot decay before noise injection further stabilizes denoising and improves success rates (+1.4%) by encouraging exploration of nearby actions (see Fig. 6).
Action chunking, which predicts short future sequences, effectively balances responsiveness and temporal consistency. Longer horizons initially enhance smoothness but eventually reduce adaptability due to accumulated open-loop errors, with 10–20 steps yielding the best trade-off. Finally, quantile-based normalization outperforms standard mean–std scaling by handling heavy-tailed action distributions, leading to more stable and robust performance across environments (see Tab. 8).
4.5 Case Analysis
Among all evaluated tasks, the Select Painting subtask in VLABench is where our $\mathcal{E}_{0}$ performs weakest (see Tab. 1). To investigate the cause, we analyzed all simulated trajectories and compared them with $\pi_{0.5}$, the strongest baseline on this task. We found that most models fail to capture the task semantics: instead of selecting the painting matching the textual description, they tend to press the middle button by default, ignoring left or right options. This behavior reduces the task to a random guess with an expected success rate of $\sim$33%.
As shown in Fig. 7, $\mathcal{E}_{0}$ exhibits more meaningful interaction by attempting to align visual and textual cues, occasionally moving toward the correct button, though sometimes failing due to limited painting-related data and partial catastrophic forgetting during fine-tuning. Notably, $\mathcal{E}_{0}$ produces smoother, more stable motions than $\pi_{0.5}$, which often acts erratically despite achieving higher success “by chance.” We also observed inconsistent success labeling in a few cases, likely due to minor issues in the benchmark code. Overall, $\mathcal{E}_{0}$ maintains strong control quality, and future work will focus on improving its visual–language grounding to enhance task understanding. Further details are provided in the section F of the appendix.
5 Conclusion
In this work, we revisited action modeling in Vision–Language–Action systems from a structural perspective. We showed that the multi-peaked nature of action distributions, the token-based reasoning of pretrained VLM/VLA backbones, and the effective finite resolution of real-world control jointly motivate discrete action modeling. Based on these insights, we proposed $\mathcal{E}_{0}$, a Tweedie discrete diffusion framework for principled and fine-grained action generation. Our results demonstrate that discrete diffusion provides a robust and generalizable foundation for action modeling in VLA systems.
Impact Statement
This work aims to advance Vision–Language–Action models for robotic manipulation by improving the robustness and generalization of action generation through a discrete diffusion-based method. As a methodological contribution, our approach is intended to support research on learning-based robotic control rather than direct deployment in safety-critical or autonomous decision-making systems.
Potential societal impacts primarily arise in real-world robotic applications, where improperly generated actions could lead to unsafe behavior or unintended interactions with the physical environment. These risks are not unique to our method and are common to learning-based robotic control in general. In practice, such risks can be mitigated through standard safeguards, including deployment in controlled environments, human supervision, simulation-based validation, and existing physical safety mechanisms on robotic platforms. We do not anticipate new ethical concerns beyond those already well established in the field of robot learning.
Appendix A Implementation and Model Details
A.1 Architecture and Pipeline
During inference, we initialize the action sequence from Gaussian noise and perform a multi-step iterative refinement procedure. At the first iteration, the observation $o_{t}$ is encoded once by the model to construct a key–value cache $\mathit{KV}(o_{t})$ for cross-attention. This cache is reused across all subsequent refinement steps, such that only the action tokens are recomputed.
We discretize a continuous refinement variable $\tau\in[1,0]$ into $N$ (with $N=10$ in our experiments) steps using a uniform schedule, i.e., $\tau_{i}=1-\frac{i-1}{N}$ for $i=1,\ldots,N$. At each iteration $i$, the current noisy action representation $\tilde{A}_{t}^{\tau_{i}}$ is passed into the model together with the fixed cache $\mathit{KV}(o_{t})$. The model produces logits $v_{\theta}(\tilde{A}_{t}^{\tau_{i}},\mathit{KV}(o_{t}))$, which define a categorical distribution over discrete action tokens:
$p_{\theta}(A_{t}\mid\tilde{A}_{t}^{\tau_{i}},o_{t})=\mathrm{Softmax}\!\left(v_{\theta}(\tilde{A}_{t}^{\tau_{i}},\mathit{KV}(o_{t}))\right).$ (10)
The predicted categorical distribution is decoded into discrete token indices via $\arg\max$ decoding, and each token is converted into a one-hot vector in a $B$-dimensional discrete action space:
$\mathrm{onehot}(k)\in\{0,1\}^{B},\quad\mathrm{onehot}(k)_{j}=\begin{cases}1,&j=k,\\ 0,&j\neq k.\end{cases}$ (11)
The resulting one-hot vectors form the intermediate action representation $\hat{A}_{t}^{(i)}$.
To obtain the input for the next refinement step, we apply an implicit noise annealing scheme controlled by $\tau_{i+1}$:
$\tilde{A}_{t}^{\tau_{i+1}}=\tau_{i+1}\,\varepsilon+(1-\tau_{i+1})\,\hat{A}_{t}^{(i)},\quad\varepsilon\sim\mathcal{N}(0,I),$ (12)
where $\varepsilon$ is fixed-variance isotropic Gaussian noise. As $\tau$ decreases monotonically, the signal-to-noise ratio increases, yielding progressively more deterministic action representations without introducing an explicit variance schedule.
After $N$ refinement steps, the final predicted token indices $\{k_{h}\}_{h=1}^{H}$ are deterministically mapped back into continuous actions by inverting the discretization:
$a_{t,h}=\mathrm{Detokenize}(k_{h}),\quad h=1,\ldots,H.$ (13)
In practice, following (Black et al., 2024; Intelligence et al., 2025), we precompute per-dimension action statistics and normalize all action values into the $[0,1]$ interval prior to discretization. This preprocessing enables efficient quantile-based discretization, where actions can be easily encoded into discrete tokens and inverse-mapped back to the original continuous action space through the corresponding rescaling operation. Collecting across the horizon yields the reconstructed action chunk:
$A_{t}=[a_{t,1},a_{t,2},\ldots,a_{t,H}].$ (14)
A.2 Spherical Perspective Generalization
To improve robustness against dynamic camera perturbations, we introduce a spherical warping augmentation and a corresponding relative spherical embedding. This design is motivated by a key limitation in current robot learning setups: Most policies are trained and validated under fixed camera viewpoints, which leads to a strong dependency on camera placement. As a result, when the camera position or orientation is altered, the inference accuracy of the policy degrades significantly. By explicitly augmenting training with spherical viewpoint perturbations and embedding the relative camera offsets into the representation space, our approach mitigates this reliance on fixed viewpoints and enables more robust generalization under dynamic or shifted camera configurations.
Spherical Warping. Given an RGB image $I\in\mathbb{R}^{H\times W\times 3}$ and camera intrinsics $(f_{x},f_{y},c_{x},c_{y})$, each pixel $(u,v)$ is first unprojected to a 3D point at fixed depth $d_{\mathrm{center}}$:
$X=\frac{(u-c_{x})\,d_{\mathrm{center}}}{f_{x}},\quad Y=\frac{(v-c_{y})\,d_{\mathrm{center}}}{f_{y}},\quad Z=d_{\mathrm{center}}.$ (15)
We then apply spherical perturbations defined by yaw ($\Delta\theta$) and pitch ($\Delta\phi$) rotations:
$R=R_{x}(\Delta\phi)\,R_{y}(\Delta\theta),\quad\begin{bmatrix}X^{\prime}\\ Y^{\prime}\\ Z^{\prime}\end{bmatrix}=R\begin{bmatrix}X\\ Y\\ Z\end{bmatrix}$. (16)
Finally, the rotated points are projected back to the image plane:
$u^{\prime}=\frac{X^{\prime}f_{x}}{Z^{\prime}}+c_{x},\quad v^{\prime}=\frac{Y^{\prime}f_{y}}{Z^{\prime}}+c_{y},$ (17)
yielding a warped image $I^{\prime}$. This augmentation simulates dynamic camera views around the robot.
Relative Spherical Embedding. To explicitly encode the relative perturbation of each camera, we define a 3D offset vector $\delta=[d,\;\theta,\;\phi]$, where $(d,\theta,\phi)$ denote radial, horizontal, and vertical displacements. For the base camera, we use $\delta_{\mathrm{orig}}=[-d,-\theta,-\phi]$, and for the warped camera, $\delta_{\mathrm{warp}}=[d,\theta,\phi]$. These offsets are mapped into the token embedding space through a learnable projection function $f_{\mathrm{proj}}$:
$e_{\mathrm{rel}}=f_{\mathrm{proj}}(\delta).$ (18)
Let $z_{\mathrm{img}}\in\mathbb{R}^{N\times d}$ denote the image tokens extracted by the vision encoder. We integrate the relative embedding additively:
$z^{\prime}_{\mathrm{img}}=z_{\mathrm{img}}+e_{\mathrm{rel}}.$ (19)
This design allows the model to not only utilize visual features but also explicitly account for the relative camera perturbations, thereby improving cross-view consistency and robustness in action generation.
Appendix B Detailed Ablation Analysis
Effect of Discretization Bins. A key advantage of our model is its ability to flexibly discretize continuous control signals into an arbitrary number of bins, enabling a fine-grained trade-off between numerical precision and representation compactness. To evaluate this, we systematically vary the discretization resolution from 128 to 8192 bins (see Tab. 4). At low resolutions (e.g., 128 bins), the model fails to learn meaningful trajectories due to the inability of coarse quantization to accurately encode joint values. This results in erratic and unstable robot motions. As the bin count increases, control accuracy improves, with performance saturating around 2048 bins. At this point, the discretization error becomes negligible compared to diffusion noise, offering sufficient precision for smooth and stable control. However, further increasing the resolution to 4096 or 8192 bins yields only marginal gains or even slight degradation in performance. This is attributed to the larger token vocabulary and increased noise, which complicate optimization. Unlike conventional autoregressive VLA models limited to 256 discrete bins, our method supports significantly higher-resolution action representations, enabling superior control fidelity and precision.
Table 4: Ablation study on discretization bins in the LIBERO environments. We evaluate different different num of bins used to discretize continuous action values. Unlike autoregressive VLAs that typically fix the bin size to 256, our model supports arbitrary discretization granularity. Performance improves with higher resolution and peaks at 2048 bins, indicating that this level of precision is sufficient for accurate action representation. Bold numbers mark the best performance per column.
bins Libero-Spatial Libero-Object Libero-Goal Libero-Long Average
SR (%) $\uparrow$ SR (%) $\uparrow$ SR (%) $\uparrow$ SR (%) $\uparrow$ SR (%) $\uparrow$
128 0.0 0.0 0.0 0.0 0.00
256 87.2 93.8 75.8 45.4 75.55
512 87.8 93.0 80.2 58.8 79.95
1024 86.2 95.6 84.6 60.2 81.65
2048 90.8 95.8 84.4 66.2 84.30
4096 89.8 95.8 83.2 65.4 83.55
8192 87.0 94.4 81.4 65.2 82.00
Effect of Action Dimension. Our model supports flexible action dimensionality when discretizing continuous control vectors. To assess the influence of this flexibility, we vary the number of action dimensions $d\in\{8,16,32\}$ during training and evaluation (see Tab. 5). We observe that setting $d$ to match or slightly exceed the maximum action dimensionality in the dataset (e.g., d=8d{=}8 in LIBERO) yields the highest average success rate. When $d$ is smaller than the true action dimension, the model cannot capture all control channels, leading to missing or corrupted actions during decoding. Conversely, using a value of $d$ significantly larger than necessary introduces numerous blank or zero-padded tokens into the sequence. These redundant tokens inject unnecessary noise into the diffusion process, degrading the sharpness and precision of predicted trajectories. Therefore, for optimal performance, $d$ should be chosen to tightly align with the dataset’s action space dimensionality, ensuring both representational completeness and robustness against embedding-induced noise.
Table 5: Ablation study on action dimension size in the LIBERO environments. We evaluate different settings of the action dimension used for representing the robot control signal. Although the model architecture supports arbitrary dimensionality, setting the dimension slightly above the maximum action size in the dataset achieves the best balance between expressiveness and noise robustness. Excessively large dimensions introduce redundant padding tokens that degrade output quality. Bold numbers mark the best performance per column.
dims Libero-Spatial Libero-Object Libero-Goal Libero-Long Average
SR (%) $\uparrow$ SR (%) $\uparrow$ SR (%) $\uparrow$ SR (%) $\uparrow$ SR (%) $\uparrow$
8 90.8 95.8 84.4 66.2 84.30
16 86.8 93.4 80.8 62.6 80.90
32 85.8 96.4 79.4 58.4 80.00
Effect of One-hot Smooth Factor. Prior to injecting noise into the one-hot action representation, we apply a one-hot smooth factor to scale the vector. This smoothing operation attenuates the sharp peaks of the one-hot distribution, thereby reducing the dominance of high-confidence logits and enabling smoother transitions within the action space during diffusion. As shown in Tab. 6, lower values of the smooth factor consistently improve success rates across all LIBERO subsets. Stronger attenuation encourages the model to better explore neighboring discrete actions and prevents overfitting to rigid one-hot boundaries. Conversely, higher smooth factor values retain excessively sharp logits, which hinder denoising stability and degrade generalization performance. These findings indicate that moderately suppressing one-hot amplitude prior to noise injection enhances the robustness of policy learning under diffusion-based action modeling.
Table 6: Ablation study on one-hot decay in the LIBERO environments. We examine the influence of the decay coefficient applied to one-hot action vectors prior to noise addition. This decay attenuates the sharp peaks of discrete logits, reducing the dominance of high-confidence tokens during diffusion. Smaller decay values lead to improved stability and higher overall success rates. Bold numbers mark the best performance per column.
decay Libero-Spatial Libero-Object Libero-Goal Libero-Long Average
SR (%) $\uparrow$ SR (%) $\uparrow$ SR (%) $\uparrow$ SR (%) $\uparrow$ SR (%) $\uparrow$
0.1 90.8 95.8 84.4 66.2 84.30
0.5 93.2 95.8 83.6 63.2 83.95
1.0 89.6 95.4 84.2 62.4 82.90
Effect of Action Chunk. We adopt an action chunking strategy, wherein the policy outputs a short sequence of future actions conditioned on the current observation. This approach has shown particular effectiveness in real-world imitation learning and VLA models (e.g., (Chi et al., 2023; Liu et al., 2024; Black et al., 2024)). By predicting multiple future actions simultaneously, the policy captures implicit temporal context, thereby mitigating non-Markovian ambiguities such as pauses, delayed responses, or lingering motions—without requiring additional historical inputs that may introduce causal confusion.
Overlapping action chunks enable temporal averaging across consecutive predictions, yielding smoother and more stable control signals. Additionally, this method allows the model to operate at a lower inference frequency while generating high-frequency motor commands—a key advantage for high-precision manipulation under constrained computational budgets. Furthermore, predicting a sequence of look-ahead actions effectively compensates for inference latency during real-world deployment, promoting consistent system dynamics and responsive robot control.
As shown in Tab. 7, increasing the action horizon initially improves performance by enhancing temporal coherence and execution smoothness. However, overly long prediction horizons degrade performance across all evaluated environments, as they reduce reactivity to environmental changes and increase compounding errors from open-loop execution. In contrast, shorter horizons (10–20 steps) provide an optimal balance between short-term responsiveness and long-term planning stability, resulting in the highest overall success rates. These findings validate our choice of a short-horizon action chunking strategy for real-world robotic deployment, where responsiveness and motion stability are equally critical.
Table 7: Ablation study on the number of action horizons in the LIBERO environments. We analyze how the length of the predicted action chunk affects the success rate across four tasks. A moderate horizon yields the highest average success rate, while both very short and very long horizons lead to degraded performance, indicating a trade-off between reactivity and temporal consistency. Bold numbers mark the best performance per column.
Action Horizon Libero-Spatial Libero-Object Libero-Goal Libero-Long Average
SR (%) $\uparrow$ SR (%) $\uparrow$ SR (%) $\uparrow$ SR (%) $\uparrow$ SR (%) $\uparrow$
1 74.8 58.0 60.4 11.2 51.10
5 84.6 80.8 72.2 28.0 66.40
10 86.0 94.6 75.6 47.0 75.80
20 78.4 95.2 79.2 52.2 76.25
30 71.4 81.4 69.8 41.6 66.05
40 49.4 82.0 64.0 38.2 58.40
50 45.6 66.6 59.2 27.2 49.65
100 31.0 57.6 47.4 14.2 37.55
Effect of Data Normalization. We compare two normalization techniques for processing continuous action vectors within our model. Both methods apply element-wise normalization based on statistical parameters (e.g., mean and standard deviation) computed independently for each action dimension across the entire dataset.
$\tilde{x}=\frac{x-\mu}{\sigma+10^{-6}},$ (20)
where $\mu\in\mathbb{R}^{d}$ and $\sigma\in\mathbb{R}^{d}$ denote the mean and standard deviation of each action dimension, respectively. This is the standard mean–std normalization that assumes a Gaussian-like distribution.
$\tilde{x}=\frac{(x-q_{0.01})}{(q_{0.99}-q_{0.01})+10^{-6}}\times 2-1,$ (21)
where $q_{0.01}$ and $q_{0.99}$ represent the 1st and 99th percentiles of the empirical action distribution for each dimension. This quantile-based normalization scales values into [1,1][-1,1] and reduces sensitivity to outliers or heavy-tailed distributions.
Both normalization parameters $(\mu,\sigma,q_{0.01},q_{0.99})$ are computed independently along each action dimension to accurately reflect the heterogeneous range and variance of robot control signals. As shown in Tab. 8, quantile normalization offers consistently improved stability and robustness across various LIBERO environments, making it a more suitable choice for diffusion-based policy learning in settings with diverse and non-uniform control signal distributions.
Table 8: Ablation study on normalization statistics in LIBERO environments . Comparison between mean–std normalization and quantile-based normalization. Bold numbers mark the best performance per column.
Normalization Type Libero-Spatial Libero-Object Libero-Goal Libero-Long Average
SR (%) $\uparrow$ SR (%) $\uparrow$ SR (%) $\uparrow$ SR (%) $\uparrow$ SR (%) $\uparrow$
Mean–Std 13.2 0.0 9.2 7.8 7.6
Quantile 90.8 95.8 84.4 66.2 84.3
Appendix C Hardware Details
We conducted real-world experiments using a Franka Emika Research 3 robotic arm, equipped with two Intel RealSense D435i cameras. One camera was mounted on the robot’s wrist, providing a dynamic view that changes with the motion of the robotic arm. The second camera was placed laterally in the workspace to capture a stable third-person perspective. We utilized Gello (Wu et al., 2024) to control the Franka arm and collect demonstration data.
In the real-world evaluation setup (see Fig. 8), we capture the visual observations required by the model using both a side-view camera and a wrist-mounted camera on the robot. In addition, a tripod-mounted camera is positioned behind and slightly to the side of the workspace to record the entire experiment for analysis and visualization.
Figure 8: Real-world evaluation setup. Observations are captured using a side-view camera and a wrist-mounted camera, while an additional tripod-mounted camera records the scene from a rear-side perspective.
Appendix D More Details on LIBERO
To evaluate the generalization behavior of our model under diverse distribution shifts, we adopt LIBERO (Liu et al., 2023) as a comprehensive benchmark for robot manipulation. LIBERO provides 130 language-conditioned tasks grouped into four procedurally generated suites (spatial, object, goal, and libero-100), each targeting a different dimension of knowledge transfer, including spatial reasoning, object concepts, behavioral goals, and mixed declarative–procedural knowledge. These tasks are created through a scalable generation pipeline that samples human-activity-inspired templates, constructs scene layouts, and specifies PDDL-based goal predicates, resulting in highly diverse manipulation scenarios with controllable variations in object types, spatial layouts, and goal semantics.
Following the common practice in prior work (use libero-10 instead of libero-100) and consistent with the configurations adopted by most existing models (Black et al., 2024; Kim et al., 2024; Pertsch et al., 2025), we train a single unified policy across all LIBERO tasks, instead of training task-specific models. All training is performed using a single RTX PRO 6000 GPU over 30,000 steps. Each task provides 50 high-quality teleoperated demonstrations, enabling sample-efficient behavioral cloning and ensuring that performance differences primarily reflect the model’s ability to transfer and retain knowledge across tasks.
As illustrated in Fig. 9, the pick up the milk and place it in the basket task reveals clear behavioral differences among the baseline models and our proposed $\mathcal{E}_{0}$. Both $\pi_{0}$ and $\pi_{0}$-FAST frequently approached the target object with suboptimal gripper orientation, often making lateral contact with the milk carton, which destabilized the object and occasionally knocked it over during the grasping phase. Although $\pi_{0.5}$ demonstrated slightly improved task semantics, it still exhibited inconsistent end-effector alignment, leading to partial or unstable grasps.
In contrast, our $\mathcal{E}_{0}$ consistently produced a precise, well-aligned grasp, approaching the milk carton with the correct wrist rotation and a controlled closing trajectory. This allowed the policy to lift the object cleanly without disturbance and reliably place it into the basket. Such stable behavior suggests that $\mathcal{E}_{0}$ learns more accurate contact priors, smoother approach trajectories, and better-conditioned fine-motor actions compared to previous policies. The improvement also highlights $\mathcal{E}_{0}$’s enhanced robustness in visually cluttered scenes and its ability to preserve fine-grained manipulation correctness, even when trained jointly across many heterogeneous LIBERO tasks.
Although the LIBERO benchmark has already been pushed to very high success rates by many recent models (see Tab. 9), it remains one of the most widely adopted and reliable platforms for stress-testing manipulation policies, particularly for evaluating robustness, motion stability, and semantic correctness under diverse visual and task variations. Despite the overall high performance saturation, our $\mathcal{E}_{0}$ still exhibits clear qualitative and quantitative advantages, demonstrating smoother behaviors, better grasp stability, and more precise action execution than strong baselines.
Figure 9: Comparison on the LIBERO benchmark. In the task pick up the milk and place it in the basket, our $\mathcal{E}_{0}$ successfully grasped the object with the correct posture and completed the task without knocking it over.
Table 9: LIBERO task performance (success rate, SR) results. Evaluation across four subsets—Spatial, Object, Goal, and Long-horizon. Diffusion-based and autoregressive VLAs are compared, showing that large-scale generalist models significantly improve robustness and our proposed $\mathcal{E}_{0}$ achieves the highest average SR, demonstrating strong cross-task generalization and stability. Bold numbers denote the best results per column.
Model Libero-Spatial Libero-Object Libero-Goal Libero-Long Average
SR (%) $\uparrow$ SR (%) $\uparrow$ SR (%) $\uparrow$ SR (%) $\uparrow$ SR (%) $\uparrow$
Diffusion Policy (Chi et al., 2023) 78.3 92.5 68.3 50.5 72.4
MDT (Reuss et al., 2024) 78.5 87.5 73.5 64.8 76.1
OpenVLA (Kim et al., 2024) 84.7 88.4 79.2 53.7 76.5
Octo (Mees et al., 2024) 78.9 85.7 84.6 51.1 75.1
Dita (Hou et al., 2025) 84.2 96.3 85.4 63.8 82.4
TraceVLA (Zheng et al., 2024) 84.6 85.2 75.1 54.1 74.8
SpatialVLA (Qu et al., 2025) 88.2 89.9 78.6 55.5 78.1
$\pi_{0}$ FAST (Pertsch et al., 2025) 96.4 96.8 88.6 60.2 85.5
$\pi_{0}$ (Black et al., 2024) 96.8 98.8 95.8 85.2 94.2
$\pi_{0.5}$ (Intelligence et al., 2025) 95.4 98.4 97.2 89.6 95.2
$\mathcal{E}_{0}$ (ours) 97.2 99.4 95.0 92.2 96.0
Appendix E More Details on ManiSkill
To further assess the generalization capability and fine-grained manipulation performance of our model across diverse operational scenarios, we additionally conduct experiments on ManiSkill (Tao et al., 2024). ManiSkill is an advanced GPU–parallelized simulation and benchmarking framework for embodied AI, offering one of the most comprehensive suites of manipulation tasks to date. It covers 12 distinct categories, ranging from tabletop manipulation and dexterous hand control to room-scale mobile manipulation, bimanual coordination, and high-fidelity digital twin environments. Its highly efficient GPU-based simulation and rendering pipeline enables large-scale visual policy training on a single GPU with minimal memory overhead, making it a suitable platform for evaluating both generalization and visual robustness.
Following the experimental design used in RDT (Liu et al., 2024), we leverage the official demonstration dataset (5 tasks with a total of 5,000 trajectories) and retrain all baseline models under identical data conditions to ensure fair comparison. All training is performed on a single RTX PRO 6000 GPU for 30,000 steps, using the same evaluation protocol across models. This setup allows us to rigorously compare their representational capacity, generalization behavior, and fine-grained manipulation performance within a high-fidelity simulation environment.
As shown in Fig. 10, the PegInsertionSide task highlights clear behavioral differences among our model and baseline models. Both $\pi_{0}$ and $\pi_{0}$-FAST struggle to produce stable insertion trajectories: their end-effector often approaches the peg with incorrect orientation, leading to repeated collisions with the box surface instead of aligning with the hole. Although $\pi_{0.5}$ demonstrates slightly improved spatial reasoning, its execution remains unstable, showing drifting motions and failed insertion attempts.
In contrast, our $\mathcal{E}_{0}$ exhibits highly reliable and precise insertion behavior. Across the entire rollout sequence, $\mathcal{E}_{0}$ maintains a well-controlled approach trajectory, accurately aligns the peg with the hole, and executes a smooth and decisive insertion. This qualitative advantage—clear from the consistent orientation control and absence of collision-induced jitter suggests that $\mathcal{E}_{0}$ learns more stable contact, aware action patterns, and benefits significantly from our discrete diffusion–based action representation.
These visual differences are mirrored by the quantitative results in Tab. 10. Across five challenging manipulation tasks—PegInsertionSide, PickCube, StackCube, PlugCharger, and PushCube—our model achieves the highest average success rate (55.2%), highlighting its balanced capability in precision manipulation, geometry-aware reasoning, and stable interactions.
Overall, the combined qualitative and quantitative evidence demonstrates that $\mathcal{E}_{0}$ achieves state-of-the-art generalization and stability on ManiSkill, and remains robust even in tasks where many of the strongest existing policies collapse. This validates the effectiveness of our discrete diffusion design in capturing both fine motor control and long-horizon structural dependencies in complex manipulation environments.
Figure 10: Comparison on the ManiSkill benchmark. In the task PegInsertionSide, our $\mathcal{E}_{0}$ achieved the best performance in this highly dexterous task. It was able to precisely align the peg with the hole and successfully insert it, while other models performed poorly.
Table 10: ManiSkill task performance (success rate, SR) results. Evaluation on five challenging manipulation tasks—PegInsertionSide, PickCube, StackCube, PlugCharger, and PushCube—from the ManiSkill benchmark. Compared with both diffusion-based and autoregressive VLAs, our proposed $\mathcal{E}_{0}$ achieves the highest average SR. While prior models often fail on precise insertion or long-horizon coordination, $\mathcal{E}_{0}$ maintains stable performance, highlighting the effectiveness of discrete diffusion-based action representation. Bold numbers denote the best results per column.
Model PegInsertionSide PickCube StackCube PlugCharger PushCube Average
SR (%) $\uparrow$ SR (%) $\uparrow$ SR (%) $\uparrow$ SR (%) $\uparrow$ SR (%) $\uparrow$ SR (%) $\uparrow$
Diffusion Policy (Chi et al., 2023) 0.0 40.0 80.0 0.0 88.0 30.2
OpenVLA (Kim et al., 2024) 0.0 8.0 80.0 0.0 8.0 4.8
Octo (Mees et al., 2024) 0.0 0.0 0.0 0.0 0.0 0.0
RDT (Liu et al., 2024) 13.2 77.2 74.0 1.2 100.0 53.6
$\pi_{0}$ FAST (Pertsch et al., 2025) 0.0 80.0 52.0 0.0 92.0 44.8
$\pi_{0}$ (Black et al., 2024) 4.0 60.0 48.0 0.0 100.0 42.4
$\pi_{0.5}$ (Intelligence et al., 2025) 8.0 56.0 56.0 4.0 92.0 43.2
$\mathcal{E}_{0}$ (ours) 24.0 76.0 72.0 4.0 100.0 55.2
Appendix F More Details on VLABench
To further assess the robustness, semantic understanding, and generalization capability of our model across diverse language-conditioned manipulation scenarios, we additionally evaluate it on VLABench (Zhang et al., 2024). VLABench offers a broad and heterogeneous collection of tasks that span various object categories, spatial arrangements, and visually complex environments. Each task is paired with natural language instructions that require accurate semantic grounding, attribute-level understanding, and precise alignment between linguistic cues and visual observations. The benchmark’s extensive variability in object layouts, appearance conditions, and linguistic expressions makes it a rigorous platform for evaluating multimodal reasoning and visuomotor robustness.
Following common practice in prior work and to ensure fair comparison, we train a single unified model rather than training separate models for individual tasks. We adopt the official VLABench dataset and retrain all baseline methods under identical data and computational settings. All experiments are conducted on a single RTX PRO 6000 GPU for 30,000 training steps, ensuring consistent training dynamics across models.
As shown in Tab. 11, our model achieves the highest overall success rate on VLABench, outperforming a variety of baseline approaches across multiple language-guided manipulation tasks. These results indicate that our discrete-diffusion action representation provides stronger semantic grounding, more stable visuomotor behavior, and improved generalization across diverse linguistic and visual conditions.
In accordance with the official VLABench protocol, we conduct mixed training using the full dataset released by the benchmark. A noteworthy aspect of this dataset is that it contains a large number of trajectories unrelated to the evaluation tasks. This increases the difficulty of generalization: models lacking sufficient robustness may overfit to these irrelevant instruction–action pairs, ultimately reducing performance on the actual evaluation tasks. As a result, VLABench serves as a particularly challenging benchmark for evaluating semantic alignment, instruction following, and cross-task transfer.
While the main text has already provided a detailed discussion of the Select Poker task, we do not repeat that analysis here. Instead, we further revisit the Select Painting task, for which a partial discussion has already been included in the main paper. As shown in Fig. 11, we provide additional examples illustrating that some predicted “failures” are in fact misjudged cases. In these instances, our model successfully identifies and selects the correct painting, yet the simulation environment still labels the outcome as failure. Such discrepancies are likely caused by limitations in the evaluation logic of the simulated environment rather than errors made by the policy itself. These observations highlight that certain failure cases are attributable to annotation or environmental inconsistencies, rather than deficiencies in the model’s reasoning or visuomotor execution.
In addition to reporting the success rate (Tab. 11), we also include the process score (Tab. 12), a complementary metric that measures whether the policy executes the correct intermediate reasoning steps even when the final prediction is annotated as failure. Our $\mathcal{E}_{0}$ achieves the highest average PS across all five evaluation tasks, indicating more accurate step-by-step action grounding and stronger adherence to instruction semantics throughout the execution sequence. Together, these results highlight the robustness, compositional understanding, and semantic reliability of our model, even in the presence of ambiguous labels and visually confounding selection tasks.
Figure 11: More comparison on the VLABench select painting task. This figure extends the results presented in Fig. 7, showcasing additional examples and outcomes.
Table 11: VLABench task performance (success rate, SR) results. Evaluation on five language-conditioned visual reasoning tasks from VLABench, including Select Toy, Select Fruit, Select Painting, Select Poker, and Select Mahjong. These tasks require fine-grained visual grounding and action understanding under compositional instructions. Compared to baseline models, our proposed $\mathcal{E}_{0}$ achieves the highest average success rate, showing stronger instruction following and cross-domain transferability. Bold numbers denote the best results in each column.
Model Select Toy Select Fruit Select Painting Select Poker Select Mahjong Average
SR (%) $\uparrow$ SR (%) $\uparrow$ SR (%) $\uparrow$ SR (%) $\uparrow$ SR (%) $\uparrow$ SR (%) $\uparrow$
$\pi_{0}$ (Black et al., 2024) 54.0 48.0 16.0 6.0 6.98 26.20
$\pi_{0}$ FAST (Pertsch et al., 2025) 46.0 42.0 26.0 30.0 20.83 32.97
$\pi_{0.5}$ (Intelligence et al., 2025) 24.0 18.0 36.0 20.0 6.52 20.90
$\mathcal{E}_{0}$ (ours) 54.0 34.0 12.0 72.0 18.75 38.15
Table 12: VLABench task performance (process score, PS) results. We report process-level performance across five language-conditioned selection tasks from VLABench (Zhang et al., 2024). While baselines such as $\pi_{0}$ (Black et al., 2024), $\pi_{0}$ FAST (Pertsch et al., 2025), and $\pi_{0.5}$ (Intelligence et al., 2025) demonstrate task-specific strengths, our proposed $\mathcal{E}_{0}$ attains the highest average process score, reflecting more consistent and fine-grained reasoning throughout multi-step action modeling. Bold numbers indicate the best performance in each column.
Model Select Toy Select Fruit Select Painting Select Poker Select Mahjong Average
PS $\uparrow$ PS $\uparrow$ PS $\uparrow$ PS $\uparrow$ PS $\uparrow$ PS $\uparrow$
$\pi_{0}$ (Black et al., 2024) 0.76 0.72 0.16 0.1000 0.0814 0.3643
$\pi_{0}$ FAST (Pertsch et al., 2025) 0.72 0.68 0.26 0.4200 0.3333 0.4827
$\pi_{0.5}$ (Intelligence et al., 2025) 0.52 0.49 0.36 0.2867 0.1739 0.4577
$\mathcal{E}_{0}$ (ours) 0.76 0.65 0.12 0.7733 0.2500 0.5107
Appendix G More Details on RoboTwin
To further systematically evaluate the robustness and generalization capability of our model, we conducted an additional experiment using RoboTwin (Chen et al., 2025; Mu et al., 2025) as the primary benchmark for dual-arm manipulation. RoboTwin provides 50 diverse tasks (see Fig. 18 and Fig. 19), covering a wide range of object categories, both single-arm and dual-arm scenarios, and various interaction patterns. Each task includes both clean and heavily domain-randomized trajectories, enabling controlled evaluation under both standard and challenging conditions. Compared to prior simulation benchmarks, RoboTwin introduces significant domain randomization, including clutter, lighting variations, background textures, tabletop height, and linguistic diversity. These features make it particularly well-suited for assessing the robustness of VLA-based manipulation models in cluttered, visually diverse, and previously unseen environments.
For a fair comparison and to demonstrate generalization, we did not train a separate model for each individual task. Instead, we mixed the initial data from all 50 official tasks and trained both our model and the baseline model $\pi_{0}$ under the same conditions. Training was performed using a single RTX PRO 6000 GPU over 30,000 steps. Evaluation was conducted using a unified random seed. As shown in the Tab. 13, our model achieved the highest average performance across all tasks, surpassing the baseline by 8.0%. It performed best on single-arm tasks, achieving a 13.7% higher average success rate. Even on the more demanding dual-arm tasks that require precise coordination, our model still outperformed the baseline, with a 1.3% higher average success rate.
Furthermore, Fig. 12 and Fig. 13 present four tasks from the single-arm and dual-arm settings, respectively, where our model $\mathcal{E}_{0}$ shows the largest performance deviations from the baseline $\pi_{0}$. These include the two tasks with the highest positive improvement in success rate and the two tasks with the largest negative performance gap. This comparison further reveals the strengths and limitations of our model across diverse manipulation scenarios.
In the single-arm tasks (Fig. 12), $\mathcal{E}_{0}$ achieves notable success over $\pi_{0}$ in Adjust Bottle and Place Object Scale. In Adjust Bottle, our model succeeds where $\pi_{0}$ fails, indicating stronger capability in precise pose correction and object reorientation. In Place Object Scale, $\mathcal{E}_{0}$ is able to perform accurate placement under spatial constraints, suggesting enhanced visual grounding and motor control. However, in Turn Switch and Open Laptop, $\mathcal{E}_{0}$ underperforms compared to $\pi_{0}$. Both tasks involve articulated or constrained motion (rotating a switch and opening a hinged object) where fine motor precision, torque control, or structural understanding are critical. These results suggest that while $\mathcal{E}_{0}$ generalizes well to visually diverse tasks, it may struggle with fine-grained control and interaction dynamics in tasks that require specific mechanical affordances.
In the dual-arm tasks (Fig. 13), $\mathcal{E}_{0}$ shows strong gains over the baseline in Place Burger Fries and Stack Blocks Two. Notably, the former involves cluttered environments, where our model maintains high success rates, suggesting better visual robustness and policy adaptability. However, in Handover Mic and Stack Bowls Three, $\mathcal{E}_{0}$ performs worse than $\pi_{0}$, revealing difficulties in tasks requiring complex bimanual coordination or long-horizon, sequential execution, especially when inter-step dependencies are critical.
Overall, these results highlight that while $\mathcal{E}_{0}$ demonstrates strong generalization across a wide range of tasks, particularly in visually complex environments and precision-driven single-arm manipulation, certain task categories still present performance bottlenecks. Tasks requiring fine-grained temporal coordination, long-horizon planning, or dual-arm synchronization continue to challenge the model’s current capabilities. One possible contributing factor is the nature of mixed-task training: when training on a large and diverse set of tasks simultaneously, specific skills, such as delicate manipulation or tightly coupled bimanual control, may be underrepresented or diluted. This can lead to suboptimal specialization on tasks that demand precise timing or coordination. Addressing these limitations by exploring more balanced task sampling, modular policy design, or adaptive task curricula will be a central focus of our future work to further enhance robustness and skill transfer in complex manipulation settings.
Figure 12: Comparison on the RoboTwin benchmark. In the single-arm tasks, our $\mathcal{E}_{0}$ shows the largest performance deviations from $\pi_{0}$ across four tasks in total—including the two tasks with the highest positive improvement in success rate and the two tasks with the largest negative gap relative to $\pi_{0}$.
Figure 13: Comparison on the RoboTwin benchmark. In the dual-arm tasks, our $\mathcal{E}_{0}$ shows the largest performance deviations from $\pi_{0}$ across four tasks in total—including the two tasks with the highest positive improvement in success rate and the two tasks with the largest negative gap relative to $\pi_{0}$.
Table 13: RoboTwin results across all 50 tasks. Red numbers indicate improvement, while green numbers indicate decline. Compared comprehensively with $\pi_{0}$, our model $\mathcal{E}_{0}$ achieves the highest average success rate across single-arm, dual-arm, and overall task categories.
Task (Left) $\pi_{0}$ SR (%) $\mathcal{E}_{0}$ SR (%) Diff Task (Right) $\pi_{0}$ SR (%) $\mathcal{E}_{0}$ SR (%) Diff
Single-Arm Tasks stamp seal 27.0 40.0 +13.0
adjust bottle 41.0 97.0 +56.0 turn switch 39.0 20.0 -19.0
beat block hammer 66.0 73.0 +7.0 Dual-Arm Tasks
click alarmclock 44.0 37.0 -7.0 blocks ranking rgb 17.0 34.0 +17.0
click bell 28.0 7.0 -21.0 blocks ranking size 7.0 11.0 +4.0
dump bin bigbin 58.0 54.0 -4.0 grab roller 86.0 94.0 +8.0
move can pot 41.0 44.0 +3.0 handover block 9.0 0.0 -9.0
move pillbottle pad 24.0 49.0 +25.0 handover mic 92.0 25.0 -67.0
move playingcard away 54.0 92.0 +38.0 hanging mug 12.0 3.0 -9.0
move stapler pad 2.0 16.0 +14.0 lift pot 32.0 40.0 +8.0
open laptop 54.0 36.0 -18.0 pick diverse bottles 36.0 42.0 +6.0
open microwave 50.0 69.0 +19.0 pick dual bottles 52.0 42.0 -7.0
place a2b left 22.0 57.0 +35.0 place bread basket 35.0 60.0 +25.0
place a2b right 14.0 48.0 +34.0 place bread skillet 34.0 47.0 +13.0
place container plate 85.0 96.0 +11.0 place burger fries 52.0 84.0 +32.0
place empty cup 51.0 88.0 +37.0 place can basket 40.0 52.0 +12.0
place fan 27.0 16.0 -11.0 place cans plasticbox 18.0 47.0 +29.0
place mouse pad 6.0 39.0 +33.0 place dual shoes 15.0 13.0 -2.0
place object scale 15.0 58.0 +43.0 place object basket 55.0 55.0 0.0
place object stand 58.0 83.0 +25.0 put bottles dustbin 31.0 4.0 -27.0
place phone stand 25.0 39.0 +14.0 put object cabinet 18.0 30.0 +12.0
place shoe 47.0 53.0 +6.0 scan object 32.0 33.0 +1.0
press stapler 52.0 91.0 +39.0 stack blocks three 17.0 10.0 -7.0
rotate qrcode 35.0 29.0 -6.0 stack blocks two 48.0 83.0 +35.0
shake bottle 92.0 95.0 +3.0 stack bowls three 58.0 28.0 -30.0
shake bottle horizontally 95.0 97.0 +2.0 stack bowls two 92.0 78.0 -14.0
Average on Single-Arm Tasks (27 tasks) 42.7 ($\pi_{0}$) $\rightarrow$ 56.4 ($\mathcal{E}_{0}$) (+13.7)
Average on Dual-Arm Tasks (23 tasks) 38.6 ($\pi_{0}$) $\rightarrow$ 39.9 ($\mathcal{E}_{0}$) (+1.3)
Average (50 tasks) 40.8 ($\pi_{0}$) $\rightarrow$ 48.8 ($\mathcal{E}_{0}$) (+8.0)
Appendix H More Details on Real-World Experiments
To demonstrate the real-world transferability, robust generalization, and fine-grained control capabilities of our model, we conducted a series of eight diverse real-world tasks using the Franka robotic arm. These tasks included five short-horizon tasks (pick a cube, press a button, stack cubes, pull a drawer, and close a door), as well as three long-horizon tasks (pick up two cubes consecutively, pull open a drawer and place a cube inside, and place a dish and close the door).
For the short-horizon tasks, we collected 50 trajectories per task and trained a unified model across all tasks. For the long-horizon tasks, we collected 80 trajectories per task and trained a separate unified model. All models were trained under identical settings on a single NVIDIA RTX 6000 GPU for 30,000 steps. Detailed comparisons with baseline models are provided and analyzed in the main paper (see Tab. 2), and thus are not repeated here.
To further evaluate the model’s generalization in real-world scenarios, we designed several qualitative tests. As illustrated in Fig. 16, the model successfully recognizes and localizes red cubes placed at varying positions on the table in the pick block task. In the close door task, it adapts to different initial opening angles of the microwave door. In the press button task, it accurately identifies small circular buttons at varying positions.
A particularly notable generalization test was performed in the stacking cubes task. During data collection, the green cube was consistently placed to the left of the red cube. However, during testing, we reversed their relative positions while keeping the instruction fixed: “Pick up the red cube and stack it on the green cube.” Remarkably, the model correctly interpreted the instruction, accurately identified the new target positions, and demonstrated strong generalization to scenarios not seen during training.
For the long-horizon tasks (e.g., pick twice), the model similarly demonstrated robust generalization. As shown in Fig. 14, despite variations in the placement distance, positions of the colored cubes and dishes, and the addition of distractor cubes with new colors, the model was able to follow task instructions with high accuracy. Notable exceptions include Fig. 14 (Row 4), where the green dish was placed near the robot’s physical limit, resulting in marginal placement, and Row 5, where the picking order of cubes was incorrect. Nevertheless, the model successfully located the green dish, intentionally placed in a corner, and placed both cubes into it.
Furthermore, Fig. 15 presents a controlled experiment validating the model’s ability to make real-time decisions based on observations. In two identical trials, we introduced manual interference in one run by shifting the cube after the robot had initially targeted it. The robot responded appropriately by re-localizing the cube and completing the task successfully. This showcases the model’s ability to reason beyond simple trajectory mimicry and adapt dynamically to environmental changes.
Figure 14: Comparison of keyframes in the real-world pick twice task under unseen scenarios. During evaluation, we tested the model across various unseen scenarios. In most cases, the model successfully completed the task. In a few cases, task quality is slightly compromised but still acceptable—for example, placement deviations caused by an oversized green dish, or changes in the order of object picking due to color variations.
Figure 15: Comparison of keyframes in real-world pick twice task with and without human intervention. We introduce human perturbation by manually shifting the target cube to disrupt the model’s original plan. After the interruption, the model is able to promptly adapt and replan its actions.
Figure 16: Comparison of keyframes in real-world short-horizon tasks.
Appendix I More Real-World Generalization Experiments
To further assess the model’s capacity for real-world generalization, we conducted an additional long-horizon pick twice task experiment in highly cluttered environments, as illustrated in Fig. 17. The experimental setup involved a table scattered with various toy vegetable models. During data collection, object placement was randomized across scenes without any memory restoration between episodes. In total, we collected 120 diverse and highly unstructured trajectories for training under these chaotic conditions.
Training was again performed using a single NVIDIA RTX 6000 GPU for 30,000 steps. During evaluation, we maintained the same principle of random object placement and scene variability. The results show that the model was capable of adapting to highly disordered environments, precisely interpreting and following language instructions to successfully complete the tasks. This experiment highlights the model’s strong environmental generalization and its robustness in real-world, unpredictable scenarios.
Figure 17: Qualitative results on the real-world Pick Vegetables Twice task under randomly arranged tabletop scenes. Each row shows one complete execution sequence consisting of two consecutive pick-and-place operations. Top row: carrot and cucumber. Middle row: potato and eggplant. Bottom row: pepper and corn. Across all settings, the surrounding distractor objects are placed in completely random configurations, demonstrating the robustness of our policy under visually diverse and cluttered real-world environments.
Appendix J Real-World Tasks Instructions
Pick block. Pick up the red block on the table.
Close door. Close the microwave oven door.
Press button. Press the green button.
Pull drawer. Pull out the top drawer.
Stack block. Pick up the red block and stack it on the green block.
Pick twice. Pick up the red and orange cubes in turn and place them in the green dish.
Open & put. Open the top drawer and pick up the red cube and put it in.
Put & close. Pick up the red plate, put it in the microwave, and close the door.
Pick vegetables twice 1. Pick up the red pepper and yellow corn in turn and place them in the green dish.
Pick vegetables twice 2. Pick up the yellow potato and purple eggplant in turn and place them in the green dish.
Pick vegetables twice 3. Pick up the red carrot and green cucumber in turn and place them in the green dish.
Figure 18: RoboTwin single-arm tasks. A total of 27 diverse single-arm manipulation tasks are included, covering a wide range of objects and actions.
Figure 19: RoboTwin dual-arm tasks. A total of 23 dual-arm manipulation tasks are included, covering a wide range of objects and actions. Some tasks require coordinated control between both arms, posing higher challenges for the model.
Appendix K Benefits of Discrete Action Modeling
Recent works advocate for discrete action models in robotics (Florence and Lynch, 2021; Shafiullah et al., 2022; Liang et al., 2025; Li et al., 2025; Luo et al., 2026), attributing their benefits to several key aspects. First, they mitigate the multi-peaked prediction failure inherent in continuous prediction, which mathematically forces a model to predict the conditional expectation $\hat{y}=\mathbb{E}[y|x]$, often resulting in invalid averaged actions; in contrast, discrete prediction minimizes Cross-Entropy to learn a probability mass function that captures distinct modes (see Section K.1). Second, discrete modeling preserves the backbone’s pre-trained vision and language capabilities—analogous to extending a vocabulary to new languages—while offering a potential path to inherit the scaling behavior of unified transformers for large-scale VLA research (see Secton K.2). Third, it facilitates architectural unification by keeping action generation inside the transformer with the same training objective (cross-entropy loss) used by VLMs, enabling unified modeling for images, language, and actions (see Section K.3). Fourth, discrete modeling matches the true quantized nature of real-world robot control, where hardware constraints like encoder resolution inherently discretize continuous signals (see Section K.4).
K.1 Addressing Multi-Peaked Prediction Failure
A significant body of literature has explicitly articulated why discrete action prediction often outperforms continuous action prediction (Florence and Lynch, 2021; Shafiullah et al., 2022).
A primary argument is that standard continuous prediction methods fail to model multi-peaked data effectively because they average conflicting modes, leading to invalid actions. As noted by Google Research (see https://research.google/blog/decisiveness-in-imitation-learning-for-robots/) (Florence and Lynch, 2021), this is critical in robotics. For example, in a pick-and-place task, a robot might successfully grasp an object by approaching from the left or the right. However, a continuous model trained on both strategies tends to output the average—driving the robot straight into the object. While the robot must be capable of precise continuous adjustments, it must first commit to a distinct high-level mode (left vs. right) rather than blending them.
This challenge is further addressed in works such as Behavior Transformer (BeT) (Shafiullah et al., 2022). The authors explicitly state that standard continuous regression fails due to this mode-averaging effect. To resolve this, BeT proposes discretizing the continuous action space to leverage the classification capabilities of sequence models. By using a Transformer to output a categorical distribution (via Softmax), the model can represent multiple distinct modes simultaneously, preserving the multi-peaked nature of the data.
The theoretical foundation for the “Multi-Peaked Prediction Failure” was rigorously analyzed in the Implicit Behavioral Cloning (IBC) paper (Florence et al., 2022). Although IBC utilizes Energy-Based Models rather than discrete tokens, its mathematical evidence—demonstrating that MSE regression collapses on non-unimodal data—serves as the primary justification for the field’s shift toward alternative architectures. Discrete tokens can be viewed as a “grid-based” implementation of the principles championed in the IBC paper.
Given that the “Theory of Multi-Peaked Distribution”—in particular, the necessity of avoiding the “mean” trap—is now widely accepted, providing a full proof is not our primary objective. We therefore present only a proof sketch below.
K.1.1 Definitions and Setup
Let $S$ denote the state/observation space (e.g., visual input) and $A\in\mathbb{R}^{d}$ the continuous action space.
Definition K.1 (Expert Policy).
The true data-generating distribution is denoted by $p(A|S=s)$. We assume access to a dataset of $n$ i.i.d. samples $\mathcal{D}=\{(s_{i},a_{i})\}_{i=1}^{n}\sim p(S,A)$.
Definition K.2 (Explicit Prediction).
A policy $\hat{\pi}:S\to A$ is parameterized as $f_{\theta}(s)$:
$\mathcal{L}(\theta)=\frac{1}{n}\sum_{i=1}^{n}\|f_{\theta}(s_{i})-a_{i}\|_{2}^{2}.$ (22)
Definition K.3 (Multi-Peaked).
For a fixed state $s$, the distribution $p(A|S=s)$ is multi-peaked if it consists of $k\geq 2$ distinct modes (local maxima) separated by regions of low probability density. Formally, there exist disjoint compact sets $M_{1},\dots,M_{k}\subset\mathbb{R}^{d}$ such that:
$p(A|S=s)\approx\sum_{j=1}^{k}w_{j}\cdot\delta_{M_{j}}(A),\quad\text{where }\sum_{j=1}^{k}w_{j}=1,\quad w_{j}>0.$ (23)
We assume separation between modes, i.e., $\|\mu_{i}-\mu_{j}\|_{2}>\epsilon$ for centers $\mu_{i},\mu_{j}$ of $M_{i},M_{j}$.
K.1.2 Theorem: The Failure Continuous Modeling on Multi-Peaked Distributions
Theorem K.4.
The population-optimal predictor $\hat{\pi}^{*}(s)=\arg\min_{f}\mathbb{E}_{A\sim p(A|S=s)}[\|f(s)-A\|_{2}^{2}]$ is the conditional expectation $\mathbb{E}[A|S=s]$. For a multi-peaked distribution $p(A|S=s)$ with separated modes, this mean lies in a low-density region between modes, yielding actions with high expert log-probability cost and leading to invalid control.
Proof.
We proceed in three steps: deriving the optimality, analyzing the geometric placement of the mean, and demonstrating the consequent performance degradation.
Step 1: Optimality implies Conditional Expectation
The population loss is defined as:
$L(f)=\mathbb{E}_{A\sim p(A|S=s)}[\|f(s)-A\|_{2}^{2}]=\int\|f(s)-a\|_{2}^{2}\,p(a|s)\,da.$ (24)
To find the minimizer, we take the functional derivative with respect to $f(s)$ and set it to zero:
$\frac{\partial L}{\partial f(s)}=2\int(f(s)-a)\,p(a|s)\,da=0\implies f(s)\int p(a|s)da=\int a\,p(a|s)\,da.$ (25)
Thus, the optimal predictor is:
$\hat{\pi}^{*}(s)=\mathbb{E}[A|S=s].$ (26)
The minimal loss achieved is the trace of the conditional covariance, $\text{Tr}(\text{Cov}(A|S=s))$, confirming that the loss strictly targets the first moment of the distribution.
Step 2: The Mean Lies in Low-Density Regions
Under the multi-peaked assumption, let the modes have centers $\mu_{1},\dots,\mu_{k}$ with weights $w_{j}$. The conditional mean is a convex combination:
$\mu^{*}=\mathbb{E}[A|S=s]=\sum_{j=1}^{k}w_{j}\mu_{j}.$ (27)
Since the modes are separated and $k\geq 2$, $\mu^{*}$ lies strictly inside the convex hull of the $\mu_{j}$’s. If we assume an idealized multi-peaked distribution where density is zero outside the modes ($p(a|s)=0$ for $a\notin\bigcup_{j}M_{j}$), then by convexity, the mean of disjoint sets lies in the gap between them.
$\mu^{*}\notin\bigcup_{j=1}^{k}M_{j}\implies p(\mu^{*}|s)\approx 0.$ (28)
Even with non-zero density $\eta$ between modes, $p(\mu^{*}|s)\ll\max_{a}p(a|s)$.
Step 3: Invalidity and Performance Degradation
We evaluate the policy using the Expert Cost (negative log-likelihood), which corresponds to the probability of the generated action under the expert’s distribution. For the MSE-optimal policy $\hat{\pi}^{*}(s)=\mu^{*}$:
$\text{Expert Cost}(\mu^{*}|s)=-\log p(\mu^{*}|s).$ (29)
As the density between modes $p(\mu^{*}|s)\to 0$ (e.g., in “set-valued” functions described in (Florence et al., 2022)), the cost approaches infinity:
$\lim_{p(\mu^{*}|s)\to 0}-\log p(\mu^{*}|s)=+\infty.$ (30)
In contrast, a policy that samples from the modes $\tilde{a}\sim p(A|S=s)$ has a finite expected cost equal to the entropy $H(p(A|S=s))$.
Physical Interpretation: In a robotics obstacle avoidance scenario, if $\mu_{1}$ corresponds to “Go Left” and $\mu_{2}$ corresponds to “Go Right”, the mean $\mu^{*}$ corresponds to “Go Straight”. Since “Go Straight” hits the obstacle, the optimal action is physically invalid despite being statistically optimal. ∎
K.1.3 Solution: Discrete Tokenization
Discrete approaches model $p(A|S=s)$ as a categorical distribution over discrete bins. The policy outputs a probability vector $q$ via a Softmax function:
$\hat{\pi}(s)=\arg\max_{j}q_{j},\quad q\sim\text{Softmax}(Wh(s)).$ (31)
Optimizing the Cross-Entropy (CE) loss:
$\mathcal{L}_{\text{CE}}=-\sum_{j}p_{j}\log q_{j},$ (32)
directly maximizes the log-likelihood of the modes. Unlike the continuous loss, CE does not force averaging; it allows the distribution to be multi-peaked (e.g., $P(\text{Left})=0.5,P(\text{Right})=0.5,P(\text{Straight})=0.0$), successfully avoiding the mean trap.
K.2 Preserving Capabilities and Knowledge of Pre-trained Vision-Language Models
Recent literature suggests that discrete modeling effectively preserves a backbone’s pre-trained vision and language capabilities—analogous to extending a vocabulary to new languages—thereby allowing VLA models to inherit the scaling properties of unified transformers (Liang et al., 2025; Li et al., 2025). Fully leveraging this pre-trained knowledge is critical; yet, while utilizing the same cross-entropy objective is hypothesized to minimize interference with the backbone’s original representations, a formal theoretical validation is currently missing. Given the widespread empirical adoption of this approach, we provide a proof sketch to substantiate this intuition rather than an exhaustive derivation. Specifically, we analyze the gradient dynamics with respect to the backbone’s hidden states $h$, demonstrating the superiority of discrete modeling by comparing the Gradient Boundedness and feature-space compatibility of the two loss functions.
K.2.1 Definitions
Let $f_{\theta}(x)$ denote the VLM backbone that produces a high-dimensional hidden state $h\in\mathbb{R}^{d}$ for a given context, where $\theta$ represents the pre-trained parameters.
1. Continuous Case.
We project $h$ to a continuous action $y\in\mathbb{R}^{k}$ using a linear head $W_{\mathrm{mse}}$:
$\mathcal{L}_{\mathrm{MSE}}=\frac{1}{2}\left\|W_{\mathrm{mse}}h-y_{\text{target}}\right\|_{2}^{2}.$ (33)
2. Discrete Case.
We project $h$ to logits over a vocabulary $V$ using a linear head $W_{\mathrm{ce}}$ (the token embedding matrix), followed by a Softmax:
$\mathcal{L}_{\mathrm{CE}}=-\log\left(\frac{\exp\left(w_{\text{target}}^{\top}h\right)}{\sum_{j\in V}\exp\left(w_{j}^{\top}h\right)}\right).$ (34)
K.2.2 Proof 1: Gradient Stability (The Lipschitz Constraint)
Proposition.
The gradient of the discrete loss is strictly bounded, whereas the gradient of the continuous loss is unbounded. Unbounded gradients cause large weight updates that move $\theta$ out of the pre-trained trusted region, leading to catastrophic forgetting.
Derivation.
1. Gradient of the Discrete Loss.
Let $p$ be the probability vector produced by the Softmax. The gradient with respect to the hidden state $h$ is
$\nabla_{h}\mathcal{L}_{\mathrm{CE}}=\sum_{j\in V}\left(p_{j}-\mathbb{I}[j=\text{target}]\right)w_{j}.$ (35)
Proof.
Let the logit for class $j$ be defined as
$z_{j}=w_{j}^{\top}h.$ (36)
The loss function is
$\mathcal{L}=-\log(p_{\text{target}})=-z_{\text{target}}+\log\left(\sum_{j\in V}\exp(z_{j})\right).$ (37)
We seek the gradient with respect to $h$. By the chain rule,
$\nabla_{h}\mathcal{L}=\sum_{j\in V}\frac{\partial\mathcal{L}}{\partial z_{j}}\frac{\partial z_{j}}{\partial h}.$ (38)
Step A: Derivative with Respect to Logits.
This is the standard Softmax–Cross-Entropy derivative.
  • If $j=\text{target}$:
    $\frac{\partial\mathcal{L}}{\partial z_{\text{target}}}=-1+\frac{\exp(z_{\text{target}})}{\sum_{k}\exp(z_{k})}=-1+p_{\text{target}}.$ (39)
  • If $j\neq\text{target}$:
    $\frac{\partial\mathcal{L}}{\partial z_{j}}=\frac{\exp(z_{j})}{\sum_{k}\exp(z_{k})}=p_{j}.$ (40)
These two cases can be unified using the indicator function:
$\frac{\partial\mathcal{L}}{\partial z_{j}}=p_{j}-\mathbb{I}[j=\text{target}].$ (41)
Step B: Derivative of Logits with Respect to the Hidden State.
Since $z_{j}=w_{j}^{\top}h$, we have
$\frac{\partial z_{j}}{\partial h}=w_{j}.$ (42)
Step C: Applying the Chain Rule.
Substituting the results from Steps A and B yields
$\nabla_{h}\mathcal{L}=\sum_{j\in V}\left(p_{j}-\mathbb{I}[j=\text{target}]\right)w_{j}.$ (43)
Since $p_{j}\in[0,1]$, the scalar error term is strictly bounded:
$\left|p_{j}-\mathbb{I}[j=\text{target}]\right|\leq 1.$ (44)
Therefore, the backward signal passed to the backbone is bounded by the norm of the embedding vectors $w_{j}$, and the gradient cannot explode regardless of model error.
2. Gradient of the Continuous Loss.
For the regression objective,
$\nabla_{h}\mathcal{L}_{\mathrm{continuous}}=W_{\mathrm{continuous}}^{\top}\left(W_{\mathrm{continuous}}h-y_{\text{target}}\right).$ (45)
The residual vector
$r=W_{\mathrm{continuous}}h-y_{\text{target}}$ (46)
is unbounded. Under domain shift or mismatched scaling between the latent space and target actions, $\|r\|$ may grow arbitrarily large.
3. Conclusion.
$\|\nabla_{h}\mathcal{L}_{\mathrm{continuous}}\|\propto\|\text{Error}\|\quad\text{vs.}\quad\|\nabla_{h}\mathcal{L}_{\mathrm{CE}}\|\leq C.$ (47)
Large gradients in the continuous case propagate into $\theta$, causing large parameter updates. In non-convex optimization, such updates move the model outside the basin of attraction formed during pre-training. In contrast, the CE loss preserves gradient scales consistent with pre-training dynamics, maintaining VLM capabilities.
K.2.3 Proof 2: Feature Space Distortion (Scale Invariance Conflict)
Proposition.
Continuous prediction imposes constraints on the magnitude of $h$ that conflict with Transformer normalization layers, whereas discrete classification is scale-invariant.
Derivation.
1. Transformer Dynamics.
Transformers rely heavily on Layer Normalization:
$\mathrm{LN}(h)=\frac{h-\mu}{\sigma}.$ (48)
In classification, predictions depend primarily on the angle (cosine similarity) between $h$ and token embeddings $w$. Scaling $h$ by a constant $c$ affects confidence but not the decision:
$\arg\max(W(ch))=\arg\max(Wh).$ (49)
Thus, semantic information is encoded mainly in the direction of $h$.
2. Conflict in the Continuous Objective.
For continuous modeling,
$y=W_{\mathrm{continuous}}h.$ (50)
To produce large-magnitude actions, the backbone must increase $\|h\|$:
$\|y\|\leq\|W_{\mathrm{continuous}}\|\cdot\|h\|.$ (51)
Minimizing $\mathcal{L}_{\mathrm{continuous}}$ therefore induces gradients that explicitly manipulate the norm of $h$ to match physical units, such as velocity or force.
3. Orthogonality of Objectives.
Let $g_{\text{text}}$ denote the gradient induced by pre-training objectives, which encourage angular separability of semantic features. The continuous-version gradient $g_{\text{robot}}$ contains a magnitude-altering component.
If this magnitude requirement conflicts with LayerNorm or the natural feature scale, the optimizer must substantially modify the internal weights $\theta$:
$\langle g_{\text{text}},g_{\text{robot}}^{\text{magnitude}}\rangle\approx 0.$ (52)
This orthogonal conflict distorts semantic feature clusters. The discrete objective $\mathcal{L}_{\mathrm{CE}}$, being scale-invariant, updates only feature direction and thus avoids this conflict.
K.3 Architectural Unification and Scaling via Discrete Action Spaces
Extensive literature advocates for unified architectures across diverse modalities, converging on the consensus that such alignment is crucial for fully exploiting scaling laws (Liang et al., 2025; Li et al., 2025; Luo et al., 2026). Given that pre-trained Vision-Language Models (VLMs)—and particularly Large Language Models (LLMs)—fundamentally operate on discrete tokens, adopting discrete representations for actions facilitates seamless architectural unification. These approaches ensures that action generation occurs within the transformer using the same Cross-Entropy training objective as the backbone, effectively treating images, language, and actions as a unified sequence. As this unification represents a foundational design paradigm widely validated in current research, we proceed based on this architectural premise without deriving a separate formal proof for its utility.
K.4 The Intrinsic Quantization of Physical Hardware
Discrete diffusion aligns with the intrinsic nature of robotic control, where hardware constraints—such as encoder resolution, control frequency, and actuation latency—inevitably discretize continuous signals. Unlike continuous approaches that assume infinite precision, our discrete modeling framework is adaptable, capable of matching the varying resolution requirements of diverse physical systems. Empirically, we observe that while control accuracy initially improves as the number of discrete bins increases, it eventually reaches a saturation point. Beyond this threshold, increasing bin density yields no further gains and may even degrade performance (see Fig. 6(a)). This phenomenon indicates that for any given robot, there exists an optimal quantization granularity dictated by practical physical dynamics rather than theoretical capacity. Crucially, this effective optimal resolution often diverges from the nominal resolution stated in manufacturer specifications, suggesting that usable precision is limited by real-world factors (e.g., noise or mechanical backlash) rather than just sensor bit-depth.
Appendix L Benefits of Tweedie Discrete Diffusion
Finally, we establish the theoretical superiority of Tweedie Discrete Diffusion Models over traditional discrete frameworks. We focus our comparison on mask-based discrete diffusion, which currently represents the state-of-the-art in this domain (Google DeepMind, 2025; Nie et al., 2025; Ye et al., ; Wu et al., 2025).
Our analysis demonstrates that Tweedie Discrete Diffusion exhibits robust convergence properties. Specifically, the minimization of the training objective (Cross-Entropy, CE) guarantees the optimization of the underlying generative dynamics (Fisher divergence). In contrast, mask-based models—owing to the absence of a Gaussian forward process—lack these bidirectional bounds. Consequently, a reduction in CE loss does not strictly imply a reduction in Fisher divergence, leading to a potential misalignment between the training objective and the generative dynamics.
L.1 Fisher Equivalence Theorem for Tweedie Discrete Diffusion
L.1.1 Theorem Statement and Physical Interpretation
Theorem L.1 (Equivalence between Excess Risk and Fisher Divergence).
Benefiting from the adoption of a Gaussian forward process, Tweedie Discrete Diffusion satisfies the following fundamental equivalence:
$\mathrm{ExcessCE}(\theta)\asymp D_{F}\!\left(p_{t}\,\|\,p_{t,\theta}\right),$ (53)
where $\mathrm{ExcessCE}(\theta)$ denotes the excess cross-entropy risk and $D_{F}(p_{t}\|p_{t,\theta})$ denotes the Fisher divergence (score matching error). The equivalence constants are explicit and depend only on the signal-to-noise ratio $\sigma_{t}^{2}/\alpha_{t}$ and a probability margin constant $\delta(\sigma_{t})$, and are independent of the number of classes $K$.
Physical Meaning and Convergence Analysis.
The above equivalence implies that during training, whenever the cross-entropy objective decreases by $\varepsilon$, the Fisher divergence necessarily decreases by $\Theta(\varepsilon)$, and vice versa. Consequently:
  • Tweedie Discrete Diffusion enjoys strong convergence guarantees: optimization of the training objective (CE) directly ensures optimization of the underlying generative dynamics (score/Fisher).
  • Mask-Based Discrete Diffusion, which does not employ a Gaussian forward process, lacks such bidirectional bounds. As a result, a reduction in CE does not necessarily correspond to a genuine decrease in Fisher divergence, potentially leading to a mismatch between the training objective and the generative dynamics.
L.1.2 Proof and Derivation
A. Forward Process and Training Objective.
In Tweedie Discrete Diffusion, the forward noising process is defined as a continuous Gaussian channel:
$x_{t}=\sqrt{\alpha_{t}}\,x_{0}+\sigma_{t}\epsilon,\qquad\epsilon\sim\mathcal{N}(0,I).$ (54)
This formulation differs fundamentally from classical discrete diffusion models based on Markov transition matrices. The model predicts the original category $x_{0}\in\{e_{1},\dots,e_{K}\}$ via a softmax distribution $q_{\theta}(k\mid x_{t})$, and is trained using the cross-entropy risk
$R(\theta)=\mathbb{E}\big[-\log q_{\theta}(x_{0}\mid x_{t})\big].$ (55)
B. Bayes Optimality and Excess Risk.
Let $\boldsymbol{\pi}(x_{t})$ denote the true posterior distribution, where
$\pi_{k}(x_{t})=\mathbb{P}(x_{0}=e_{k}\mid x_{t}).$ (56)
By standard results in statistical learning theory, the Bayes-optimal predictor
$q^{*}(\cdot\mid x_{t})=\boldsymbol{\pi}(x_{t})$ (57)
minimizes the expected risk, yielding
$R^{*}=\min_{\theta}R(\theta)=\mathbb{E}\big[H(\boldsymbol{\pi}(\cdot\mid x_{t}))\big].$ (58)
The excess cross-entropy risk is therefore given by
$\mathrm{ExcessCE}(\theta):=R(\theta)-R^{*}=\mathbb{E}_{x_{t}}\!\left[\mathrm{KL}\big(\boldsymbol{\pi}(\cdot\mid x_{t})\,\|\,q_{\theta}(\cdot\mid x_{t})\big)\right].$ (59)
C. Core Bridge: Tweedie’s Identity.
Since the forward process is Gaussian, Tweedie’s identity relates the posterior expectation to the score of the marginal distribution:
$\nabla_{x_{t}}\log p_{t}(x_{t})=\frac{\sqrt{\alpha_{t}}}{\sigma_{t}^{2}}\big(\mathbb{E}[x_{0}\mid x_{t}]-x_{t}\big)=\frac{\sqrt{\alpha_{t}}}{\sigma_{t}^{2}}\big(\boldsymbol{\pi}(x_{t})-x_{t}\big).$ (60)
This Tweedie’s formula is proved as follows.
Proof.
We first consider the standard additive Gaussian case $x_{t}=x_{0}+\sigma\epsilon$.
$\mathbb{E}[x_{0}|x_{t}]$ $=\int_{-\infty}^{\infty}x_{0}p(x_{0}|x_{t})dx_{0}$ (61)
$=\int_{-\infty}^{\infty}x_{0}\frac{p(x_{t}|x_{0})p(x_{0})}{p(x_{t})}dx_{0}$
$=\frac{\int_{-\infty}^{\infty}x_{0}p(x_{t}|x_{0})p(x_{0})dx_{0}}{p(x_{t})}$
$=\frac{\int_{-\infty}^{\infty}x_{0}\frac{1}{\sqrt{2\pi\sigma_{t}^{2}}}e^{-\frac{(x_{t}-x_{0})^{2}}{2\sigma_{t}^{2}}}p(x_{0})dx_{0}}{p(x_{t})}$
$=\frac{\int_{-\infty}^{\infty}\Big[(x_{0}-x_{t})\frac{1}{\sqrt{2\pi\sigma_{t}^{2}}}e^{-\frac{(x_{t}-x_{0})^{2}}{2\sigma_{t}^{2}}}p(x_{0})+x_{t}\frac{1}{\sqrt{2\pi\sigma_{t}^{2}}}e^{-\frac{(x_{t}-x_{0})^{2}}{2\sigma_{t}^{2}}}p(x_{0})\Big]dx_{0}}{p(x_{t})}$
$=\frac{\int_{-\infty}^{\infty}\Big[\sigma_{t}^{2}\frac{(x_{0}-x_{t})}{\sigma_{t}^{2}}\frac{1}{\sqrt{2\pi\sigma_{t}^{2}}}e^{-\frac{(x_{t}-x_{0})^{2}}{2\sigma_{t}^{2}}}p(x_{0})+x_{t}\frac{1}{\sqrt{2\pi\sigma_{t}^{2}}}e^{-\frac{(x_{t}-x_{0})^{2}}{2\sigma_{t}^{2}}}p(x_{0})\Big]dx_{0}}{p(x_{t})}$
$=\frac{\int_{-\infty}^{\infty}\sigma_{t}^{2}\frac{(x_{0}-x_{t})}{\sigma_{t}^{2}}\frac{1}{\sqrt{2\pi\sigma_{t}^{2}}}e^{-\frac{(x_{t}-x_{0})^{2}}{2\sigma_{t}^{2}}}p(x_{0})dx_{0}+\int_{-\infty}^{\infty}x_{t}\frac{1}{\sqrt{2\pi\sigma_{t}^{2}}}e^{-\frac{(x_{t}-x_{0})^{2}}{2\sigma_{t}^{2}}}p(x_{0})dx_{0}}{p(x_{t})}$
$=\frac{\sigma_{t}^{2}\int_{-\infty}^{\infty}\frac{d\Big[\frac{1}{\sqrt{2\pi\sigma_{t}^{2}}}e^{-\frac{(x_{t}-x_{0})^{2}}{2\sigma_{t}^{2}}}\Big]}{dx_{t}}p(x_{0})dx_{0}+\int_{-\infty}^{\infty}x_{t}\frac{1}{\sqrt{2\pi\sigma_{t}^{2}}}e^{-\frac{(x_{t}-x_{0})^{2}}{2\sigma_{t}^{2}}}p(x_{0})dx_{0}}{p(x_{t})}$
$=\frac{\sigma_{t}^{2}\int_{-\infty}^{\infty}\frac{d\Big[p(x_{t}|x_{0})\Big]}{dx_{t}}p(x_{0})dx_{0}+\int_{-\infty}^{\infty}x_{t}p(x_{t}|x_{0})p(x_{0})dx_{0}}{p(x_{t})}$
$=\frac{\sigma_{t}^{2}\frac{d}{dx_{t}}\int_{-\infty}^{\infty}p(x_{t}|x_{0})p(x_{0})dx_{0}+x_{t}\int_{-\infty}^{\infty}p(x_{t}|x_{0})p(x_{0})dx_{0}}{p(x_{t})}$
$=\frac{\sigma_{t}^{2}\frac{dp(x_{t})}{dx_{t}}+x_{t}p(x_{t})}{p(x_{t})}$
$=\sigma_{t}^{2}\frac{d}{dx_{t}}\log p(x_{t})+x_{t}$
i.e.,
$\boxed{\mathbb{E}[x_{0}|x_{t}]=x_{t}+\sigma_{t}^{2}\nabla_{x_{t}}\log p_{t}(x_{t})}.$ (62)
When the diffusion involves a scaling factor $\sqrt{\alpha_{t}}x_{0}$, we apply a variable transformation
$x_{t}=\sqrt{\alpha_{t}}\,x_{0}+\sigma_{t}\,\epsilon,$
equivalently
$x_{t}^{\prime}=\frac{x_{t}}{\sqrt{\alpha_{t}}}=x_{0}+\frac{\sigma_{t}}{\sqrt{\alpha_{t}}}\epsilon.$
Based on $x_{t}^{\prime}=\frac{x_{t}}{\sqrt{\alpha_{t}}}=x_{0}+\frac{\sigma_{t}}{\sqrt{\alpha_{t}}}\epsilon$ and Eqn. (62), we have:
$\mathbb{E}[x_{0}|x_{t}^{\prime}]$ $=x_{t}^{\prime}+\frac{\sigma_{t}^{2}}{\sqrt{\alpha_{t}}}\nabla_{x_{t}^{\prime}}\log p_{t}(x_{t}^{\prime})$ (63)
$=\frac{x_{t}}{\sqrt{\alpha_{t}}}+\frac{\sigma_{t}^{2}}{\alpha_{t}}\sqrt{\alpha_{t}}\nabla_{x_{t}}\log p_{t}(x_{t}^{\prime}),$
which yields:
$\boxed{\mathbb{E}[x_{0}|x_{t}]=\frac{1}{\sqrt{\alpha_{t}}}(x_{t}+\sigma_{t}^{2}\nabla_{x_{t}}\log p_{t}(x_{t})}.$
An analogous identity holds for the model distribution $p_{t,\theta}$. Comparing the two yields the pointwise relation
$\boldsymbol{\pi}(x_{t})-q_{\theta}(\cdot\mid x_{t})=\frac{\sigma_{t}^{2}}{\sqrt{\alpha_{t}}}\big(\nabla\log p_{t}(x_{t})-\nabla\log p_{t,\theta}(x_{t})\big).$ (64)
Taking the squared expectation under $p_{t}$ gives an exact proportionality between posterior error and Fisher divergence:
$\mathbb{E}_{p_{t}}\big\|\boldsymbol{\pi}-q_{\theta}\big\|_{2}^{2}=\frac{\sigma_{t}^{4}}{\alpha_{t}}\cdot 2\,D_{F}\!\left(p_{t}\,\|\,p_{t,\theta}\right).$ (65)
D. Establishing Bidirectional Bounds.
Using (65), we relate the excess cross-entropy risk to the Fisher divergence.
Lower bound (Pinsker’s inequality). By Pinsker’s inequality, $\mathrm{KL}(P\|Q)\geq\tfrac{1}{2}\|P-Q\|_{2}^{2}$, we obtain
$\mathrm{ExcessCE}(\theta)\geq\frac{1}{2}\mathbb{E}\big\|\boldsymbol{\pi}-q_{\theta}\big\|_{2}^{2}=\frac{\sigma_{t}^{4}}{\alpha_{t}}D_{F}\!\left(p_{t}\,\|\,p_{t,\theta}\right).$ (66)
Upper bound (strong convexity of KL). Assume that Gaussian convolution ensures posterior probabilities are bounded away from 0 and $1$, i.e., there exists $\delta(\sigma_{t})\in(0,\tfrac{1}{2}]$ such that $\pi_{k}(x_{t})\in[\delta,1-\delta]$ almost everywhere. Then the KL divergence satisfies
$\mathrm{KL}(P\|Q)\leq\frac{1}{2\delta(1-\delta)}\|P-Q\|_{2}^{2}.$ (67)
Combining this with (65) yields
$\mathrm{ExcessCE}(\theta)\leq\frac{\sigma_{t}^{4}}{\alpha_{t}\,\delta(1-\delta)}D_{F}\!\left(p_{t}\,\|\,p_{t,\theta}\right).$ (68)
Conclusion.
The excess cross-entropy risk is thus sandwiched between constant multiples of the Fisher divergence, completing the proof.